### v1

# Single-loop Projection-free and Projected Gradient-based Algorithms for Nonconvex-concave Saddle Point Problems with Bilevel Structure

Abstract
In this paper, we explore a class of constrained saddle point problems with a bilevel structure, where the upper-level objective function is nonconvex-concave and smooth subject to a strongly convex lower-level objective function. This class of problems finds wide applicability in machine learning, including robust multi-task learning, adversarial learning, and robust meta-learning. While some studies have focused on simpler formulations, e.g., when the upper-level objective function is linear in the maximization component, there remains a significant gap in developing efficient projection-free and projection-based algorithms with theoretical guarantees for more general settings. To bridge this gap, we propose efficient single-loop one-sided projection-free, and fully projection-based primal-dual methods. By leveraging regularization and nested approximation techniques, we initially devise a bilevel primal-dual one-sided projection-free algorithm, requiring $\mathcal{O}(\epsilon^{-4})$ iterations to attain an $\epsilon$-stationary point. Subsequently, we develop a bilevel primal-dual fully projected algorithm, capable of achieving an $\epsilon$-stationary solution within $\mathcal{O}(\epsilon^{-5})$ iterations. To the best of our knowledge, our proposed algorithms represent among the first methods for solving general non-bilinear saddle point problems with a bilevel structure.
1 Introduction
Bilevel optimization is an important class of optimization problems where one optimization problem is nested within another. Many learning and inference problems take this form with two levels of hierarchy. On the other hand, Saddle Point (SP), also known as min-max, optimization problems have received a lot of attention recently due to their versatile formulation covering a broad class of optimization problems, e.g., convex optimization with nonlinear conic constraints, which itself includes LP, quadratic programming (QP), quadratically constrained quadratic programming (QCQP), second-order cone programming (SOCP), and semidefinite programming (SDP) as its subclasses. From a broader perspective, in this paper, we study a saddle point problem with bilevel structure (SPB) of the following form
$\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}~{}\mathcal{L}(x,y)\triangleq\Phi(x,\theta^{*}(x),y)$ (1)
[…]
where $\Phi:\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{d}\to\mathbb{R}$ is a continuously differentiable function such that $\Phi(x,\theta,\cdot)$ is concave for any given $x\in\mathcal{X}$ and $\theta\in\mathbb{R}^{m}$; $\mathcal{X}\subset\mathbb{R}^{n}$ and $\mathcal{Y}\subset\mathbb{R}^{d}$ are convex and compact sets; moreover, $g:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ is a continuously differentiable and strongly convex in $\theta$. This problem simplifies to bilevel optimization when the upper-level objective function $\Phi$ is separable in the maximization variable and it encompasses saddle point problems when there is no lower-level problem, i.e., $\Phi$ is independent of the second argument. The problem formulated in (1) has a wide range of applications including robust multi-task learning [39], robust meta-learning [19, 24], adversarial learning [57], representation learning and hyperparameter optimization [20], multi-task deep AUC maximization, and multi-task deep partial AUC maximization [25].
1.1 Literature Review
In this section, we will begin by briefly discussing research in bilevel optimization and SP problems. Then, we will touch upon existing methods for solving SPB problems.
Bilevel Optimization. Researchers have proposed numerous algorithms for solving bilevel optimization problems since its initial introduction by Bracken and McGill in 1973 [5]. These methods include but are not limited to, constraint problem [40, 49] and unconstrained [26, 44] approaches. In recent years, there has been a growing research focus on applying gradient-based settings to solve bilevel problems, incorporating techniques such as implicit differentiation [12, 15, 17] and iterative differentiation [13, 18, 38, 47]. In the literature, several single and double-loop approaches have been developed to address unconstrained bilevel optimization problems [10, 28, 30, 32, 35, 54]. There are fewer studies focusing on constrained bilevel optimization problems, particularly those that consider a class of bilevel optimization problems with an upper-level objective function over a constraint set. Researchers have proposed a number of methods, including projection-based [15, 23], and projection-free [1, 2] algorithms for solving constrained bilevel optimization problems. Among the projection-based methods, specifically, the authors in [23] proposes a two-timescale stochastic approximation (TTSA) algorithm for solving a bilevel optimization problem. In this scenario, the upper-level problem is constrained and has a nonconvex objective function, while the lower-level problem is unconstrained and strongly convex. By employing a projection onto the set $\mathcal{X}$ at each iteration, their algorithm achieves a complexity of $\mathcal{O}(\epsilon^{-5/2})$ in this setting. In contrast, there is an increasing interest in using projection-free methods, which aim to reduce the computational cost associated with projections at each iteration through the utilization of a linear solver for solving the constrained setting. Recently, researchers in [1] study a single-loop projection-free method called inexact bilevel conditional gradient (IBCG) for solving a bilevel optimization problem with a smooth and constrained upper-level objective function and a strongly convex lower-level problem. Their method converges to an $\epsilon$-optimal solution with a complexity of $\mathcal{O}(\epsilon^{-2})$ when the upper-level objective function is non-convex. The convergence rates for both the TTSA and IBCG algorithms match those of their single-level counterparts.
Saddle Point (SP) Problem. In the context of saddle point (SP) problems, some research has focused on the convex-concave setting [6, 21, 41, 43], while there is an increasing number of studies concentrating on developing algorithms for nonconvex-concave setting [4, 36]. In the nonconvex-concave setting, primarily two types of algorithms, namely projected gradient-based algorithms and projection-free-based algorithms, are studied. The projected gradient-based algorithms can be applied in various methods, including but not limited to adding a momentum term [21, 52], accelerated gradient methods [22, 29], gradient descent-ascent [11, 55], multi-loop algorithms [37, 42]. To ensure a guaranteed convergence rate, numerous methods, including smoothing techniques [56] and acceleration methods [42], have been investigated. Many of these methods are based on solving regularized subproblems inexactly, leading to a multi-loop approach that can present implementation challenges. To overcome such complexities, there has been a growing preference for single-loop primal-dual methods. However, fewer research studies cover single-loop algorithms for the nonconvex-concave setting. In [53], the authors introduce a unified single-loop alternating gradient projection (AGP) algorithm designed for solving smooth nonconvex-(strongly) concave problems. Their algorithm can find an $\epsilon$-stationary point of the objective function, achieving a complexity of $\mathcal{O}(\epsilon^{-2})$ and $\mathcal{O}(\epsilon^{-4})$ under the nonconvex-strongly concave and nonconvex-concave settings, respectively.In the optimization literature, limitations in projection-based techniques have prompted researchers to explore novel and efficient projection-free approaches [31, 33, 34, 46, 51]. In the context of saddle point (SP) problems, a limited amount of research has been conducted on projection-free SP problems [4, 9, 16]. In particular, [4] investigates single-loop fully and one-sided projection-free algorithms designed to solve a saddle point problem with a smooth nonconvex-concave objective function. Their proposed method can identify an $\epsilon$-stationary solution within $\mathcal{O}(\epsilon^{-6})$ and $\mathcal{O}(\epsilon^{-4})$ for fully and one-sided projection-free methods, respectively.
Saddle Point Problem with Bilevel Structure (SPB). Within the existing literature, the focus on the SPB problem, as presented in Problem (1), is limited. [25] proposes two simple, single-loop, single-timescale stochastic methods with randomized block-sampling to solve the general form of Problem (1). This is done under the assumption that the upper-level objective function $\Phi(x,\theta^{*}(x),y)$ is a smooth, unconstrained, nonconvex strongly-concave function, and the lower-level objective function $g(x,\theta)$ is smooth and strongly-convex. Theoretically, they show that the algorithms converge to an $\epsilon$-stationary point with complexity $\mathcal{O}(\epsilon^{-4})$ under a general unbiased stochastic oracle model. In these approaches, Stochastic Gradient Descent (SGD) is utilized to update selected $\theta_{i}$ for their respective lower-level problems, Stochastic Gradient Ascent (SGA) is employed for updating the selected $y_{i}$, and a momentum update is applied to the primal variable $x$ based on the sampled $y_{i}$ and $\theta_{i}$. Contrastingly, the upper-level objective function in our problem is nonconvex-concave over a constraint set. Additionally, we propose a projection-free method to alleviate the computational cost associated with the application of projection-based methods. In another study, [19] employs a minimax bilevel optimization problem to formulate the task-robust meta-learning problem. They consider a minimax bilevel optimization problem of the form $\min_{x\in\mathcal{X}}\max_{i\in[n]}\Phi_{i}(x,\theta^{*}(x))$ subject to $\theta^{*}(x)\in\operatorname*{\arg\!\min}_{y_{i}\in\mathcal{Y}_{i}=\mathbb{R}^{d_{y_{i}}}}g_{i}(x,y_{i}),\forall i\in[n]$, a special case of Problem (1), assuming that the objective function $\Phi$ is linear in $y$. Their proposed algorithm, TaRo-BOBA (Task Robust Bilevel Optimization Bi-timescale Algorithm), is a single-loop, two-timescale, stochastic algorithm based on gradient descent-ascent to solve their minimax bilevel optimization problem. They assume that their upper-level problem is constraint nonconvex-concave, and the lower-level objective function is strongly-convex and unconstrained. According to theoretical studies, TaRo-BOBA converges to the first-order stationary point at a rate of $\mathcal{O}(\sqrt{n}\epsilon^{-5})$, where $n$ denotes the number of tasks. In a recent study [20], a minimax multi-objective bilevel optimization problem is explored, with applications in robust machine learning. They propose a single-loop Multi-Objective Robust Bilevel Two-timescale optimization algorithm, MORBiT, in a similar setting and convergence rate as discussed in [19]. In [24], the authors propose a single-loop, single-timescale stochastic algorithm using a stochastic estimator based on moving average (SEMA) to solve a nonconvex minimax bilevel optimization problem. Specifically, they investigate a problem where the upper-level objective function is nonconvex and strongly-concave, while the lower-level objective is considered strongly-convex. They theoretically demonstrate that their method can achieve an $\epsilon$-stationary solution within $\mathcal{O}(\epsilon^{-4})$ iterations, which yields better results compared to a similar study in [19].
There are a limited number of studies considering the SPB problem. Existing methods primarily focus on projection-based algorithms, and, to the best of our knowledge, there is no exploration of any projection-based and projection-free algorithm alternatives in this setup. Therefore, in this paper, we concentrate on developing both projection-free and projection-based methods for solving Problem (1).
1.2 Contribution
In this paper, we consider a class of constrained saddle point problems with a bilevel structure, where the upper-level objective function is nonconvex-concave and smooth over a compact and convex constraint set, subject to a strongly convex lower-level objective function. This study extends the current literature, which has primarily focused on using projection-based methods to solve nonconvex strongly-concave upper-level objective functions [24, 25], or where their upper-level objective function is linear in the dual variable [19, 20]. Our motivation stems from the need to develop efficient algorithms both projection-free and projection-based with theoretical guarantees, addressing the gap in methods for solving general non-bilinear SPB problem.
Utilizing regularization and nested approximation, we first develop a single-loop, one-sided projection-free algorithm. In the upper-level objective function, this algorithm uses a linear minimization oracle (LMO) and a regularized projection gradient ascent to find directions for primal and dual steps, respectively. Next, we develop a regularized single-loop projected gradient-based algorithm that uses a projection oracle (PO) to update the primal and dual steps of the upper-level objective function. In both algorithms, we incorporate a one-step gradient-type update to estimate the lower-level optimal solution and its derivative.
Our main theoretical guarantees for the proposed methods are summarized as follows:
  • The proposed bilevel primal-dual one-sided projection-free method achieves an $\epsilon$-stationary solution within $\mathcal{O}(\epsilon^{-4})$ iterations.
  • The proposed bilevel primal-dual fully projected method requires $\mathcal{O}(\epsilon^{-5})$ iterations to find an $\epsilon$-stationary point.
1.3 Motivating Examples
There are some applications of SPB formulation in Problem (1), including machine learning problems such as adversarial learning [57], representation learning and hyperparameter optimization [20], robust multi-task learning [39], robust meta-learning [19, 24], multi-task deep AUC maximization, and multi-task deep partial AUC maximization [25]. Next, we describe two examples in detail.
Adversarial Learning: With the advancement of technology, it is crucial to enhance the robustness and security of machine learning models against adversarial attacks. Adversarial learning aids in defense against attacks and enhances generalization by training models to withstand adversarial examples. Recently, in [57], the authors propose a bilevel form of adversarial training (AT), representing a unified bilevel optimization problem with a discriminator model in the upper-level and a generator model in the lower-level. Specifically, their proposed algorithm is interpreted as a bilevel optimization with a linearized lower-level problem. However, their formulation assumes knowledge of the data distribution, which is problematic in real-world applications.
To address this, one approach is to utilize distributionally robust optimization (DRO) to develop models resilient to variations in data distribution. Instead of relying on a specific assumed distribution, DRO considers a set of possible distributions, represented by an uncertainty set. This approach ensures that the solution performs well across a range of possible distributions, enhancing robustness to uncertainties in the data. Therefore, by combining the ideas of AT in bilevel optimization and DRO, we can present the problem generally as follows:
$\min_{x\in\mathcal{X}}\max_{P\in\mathcal{P}}\mathbb{E}_{P}[\hat{\ell}(x,\theta^{*}(x);\mathcal{D}^{tr})]$
[…]
where $\hat{\ell}$ and $\ell_{atk}$ represent the training loss function and the attack objective function, respectively. $\mathcal{X}$ denotes a convex set, and $\mathcal{P}$ signifies an uncertainty set. Furthermore, we can revisit Problem (1.3) by utilizing sample average approximation to approximate the expectation, and by linearizing $\ell_{atk}$, we can formulate it similarly to Problem (1) as:
$\min_{x\in\mathcal{X}}\max_{P\in\mathcal{P}}\frac{1}{n}\sum_{i=1}^{n}P_{i}\hat{\ell}(x,\theta^{*}(x);\mathcal{D}^{tr}_{i})$
[…]
where $P_{i}$ is the probability distribution for data point $\mathcal{D}^{tr}_{i}$ for $i\in[N]\triangleq\{1,2,…,n\}$, $\overline{\theta}$ is given, and $\mathcal{R}(\theta)$ is a regularization term such as $L_{2}$ regularization.
Robust Multi-task Learning: Consider a multi-task learning (MTL) problem where the goal is to train a model on multiple related tasks simultaneously, aiming to create specialized models for each task that perform better compared to training separate models for each task. The common approach to train the model for MTL is to minimize the average of all task losses. However, a drawback of this method is that all tasks are considered equally, while it is possible that the distribution of tasks is heterogeneous, causing a skewness in the performance of the acquired model toward certain tasks.
To overcome this problem, one approach is to minimize the worst task loss of the model, making the attained model robust against the distribution differences among tasks. This can be viewed as an SPB problem, wherein the upper level (UL), we have a saddle point problem, and in the lower level (LL), we have a minimization problem parameterized by the solution of the UL problem. This formulation has arisen in many examples. In task robust meta-learning problem [19, 24], a similar structure is utilized, where the primal variable is considered as a meta-learner, and the shared parameter is used as the task-specific learners in the adaptation phase. In the robust representation learning problem [20], a similar formulation is employed, where a shared representation network is regarded as the upper-level (UL) problem, and the lower-level (LL) problem is viewed as a per-task linear model parameterized with the solution of the UL problem, aiming to learn a shared representation and per-task models so that each task generalizes effectively and achieves robust performance across tasks. Generally, this problem can be expressed as follows:
$\min_{x\in\mathcal{X}}\max_{y\in\Delta_{N}}\sum_{i=1}^{N}y_{i}\ell_{i}(x,\theta_{i}^{*}(x);\mathcal{D}_{i}^{val})$
[…]
where $\mathcal{X}$ is a convex set (e.g. $\|x\|_{1}\leq r$), $[N]\triangleq\{1,2,…,n\}$ is the index of each learning task, and $\Delta_{N}$ is the simplex $\Delta_{N}\subset\mathbb{R}^{n}=\{y\in\mathbb{R}^{n}|y_{1}+…+y_{n}=1,y_{i}\geq 0\}$. In this setup, the lower-level and upper-level objectives correspond to a learning task $i\in[n]$ with its training and validation dataset pair $\mathcal{D}_{i}^{tr},\mathcal{D}_{i}^{val}$, respectively. In Problem (1.3), $\mathcal{R}(\theta)$ is a regularization term such as $L_{2}$ regularization, utilized to ensure the strong convexity of the lower-level problem.
An alternative formulation for robust MTL can be considered as a weight selection for MTL-DRO. Several research [8, 27, 45, 50] consider the concept of sample reweighting, which can be formulated as a bilevel optimization problem. In another study, utilizing distributionally robust optimization (DRO), [39] proposes a robust multi-task learning (MTL) problem in a single-level saddle point structure where the worst-task loss of the model is minimized at each step. Considering both ideas together, the problem can be formed as an SPB as follows:
$\min_{x\in\Delta_{N}}\max_{y\in\mathcal{U}}\sum_{i=1}^{N}y_{i}\ell_{i}(\theta^{*}(x);\mathcal{D}_{i}^{val})$
[…]
where $\Delta_{N}$ is the simplex $\Delta_{N}\subset\mathbb{R}^{n}=\{x\in\mathbb{R}^{n}|x_{1}+…+x_{n}=1,x_{i}\geq 0\}$, $\mathcal{U}$ is an uncertainty set, and $\mathcal{R}(\theta)$ denotes a regularization term for ensuring strong convexity of the lower-level problem.
2 Preliminaries
2.1 Assumptions and Definitions
In this section, we state the necessary assumptions required for the upper-level min-max objective function $\Phi$, the lower-level objective function $g$, and the constraint sets $\mathcal{X}$ and $\mathcal{Y}$. Moreover, we provide definitions essential for analyzing our proposed methods throughout the paper.
The upper-level objective function $\Phi$ is smooth and concave in $y$.
Assumption 2.1 (Upper-level min-max objective function).
$\Phi:\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{d}\to\mathbb{R}$ is a continuously differentiable function such that for any given $x,\theta$, $\Phi(x,\theta,\cdot)$ is a concave function. Moreover, $\nabla_{x}\Phi(x,\theta,y)$, $\nabla_{\theta}\Phi(x,\theta,y)$, and $\nabla_{y}\Phi(x,\theta,y)$ are Lipschitz continuous w.r.t $(x,\theta,y)$ such that for any $x,\overline{x}\in\mathbb{R}^{n}$ and $\theta,\overline{\theta}\in\mathbb{R}^{m}$, and $y,\overline{y}\in\mathbb{R}^{d}$
  1. 1.
    $\left\|\nabla_{x}\Phi(x,\theta,y)-\nabla_{x}\Phi(\overline{x},\overline{\theta},\overline{y})\right\|\leq L_{xx}^{\Phi}\left\|x-\overline{x}\right\|+L_{x\theta}^{\Phi}\left\|\theta-\overline{\theta}\right\|+L_{xy}^{\Phi}\left\|y-\overline{y}\right\|$,
  2. 2.
    $\left\|\nabla_{\theta}\Phi(x,\theta,y)-\nabla_{\theta}\Phi(\overline{x},\overline{\theta},\overline{y})\right\|\leq L_{\theta x}^{\Phi}\left\|x-\overline{x}\right\|+L_{\theta\theta}^{\Phi}\left\|\theta-\overline{\theta}\right\|+L_{\theta y}^{\Phi}\left\|y-\overline{y}\right\|$,
  3. 3.
    $\left\|\nabla_{y}\Phi(x,\theta,y)-\nabla_{y}\Phi(\overline{x},\overline{\theta},\overline{y})\right\|\leq L_{yx}^{\Phi}\left\|x-\overline{x}\right\|+L_{y\theta}^{\Phi}\left\|\theta-\overline{\theta}\right\|+L_{yy}^{\Phi}\left\|y-\overline{y}\right\|$.
Moreover, we impose the following common assumption on the lower-level objective function $g$.
Assumption 2.2 (Lower-level objective function).
$g(x,\theta)$ satisfies the following conditions:
  1. (i)
    For any given $x\in\mathcal{X},g(x,\cdot)$ is twice continuously differentiable. Moreover, $\nabla_{\theta}g(\cdot,\cdot)$ is continuously differentiable.
  2. (ii)
    For any given $x\in\mathcal{X},\nabla_{\theta}g(x,\cdot)$ is Lipschitz continuous with constant $L_{g}\geq 0$. Moreover, for any $\theta\in\mathbb{R}^{m},\nabla_{\theta}g(\cdot,\theta)$ is Lipschitz continuous with constant $C_{\theta x}^{g}\geq 0$.
  3. (iii)
    For any $x\in\mathcal{X},g(x,\cdot)$ is $\mu_{g}$-strongly convex with modulus $\mu_{g}>0$.
  4. (iv)
    For any given $x\in\mathcal{X},\nabla^{2}_{\theta x}g(x,\theta)\in\mathbb{R}^{n\times m}$ and $\nabla^{2}_{\theta\theta}g(x,\theta)$ are Lipschitz continuous w.r.t $(x,\theta)\in\mathcal{X}\times\mathbb{R}^{m}$, and with constant $L_{\theta x}^{g}\geq 0$ and $L_{\theta\theta}^{g}\geq 0$, respectively.
Remark 2.1.
Considering Assumption 2.2-(ii), we can conclude that $\left\|\nabla^{2}_{\theta x}g(x,\theta)\right\|$ is bounded with constant $C^{g}_{\theta x}$ for any $(x,\theta)\in\mathcal{X}\times\mathbb{R}^{m}$.
Assumption 2.3.
There exists $C_{\theta}^{\Phi}\geq 0$ such that $\left\|\nabla_{\theta}\Phi(x,\theta^{*}(x),y)\right\|\leq C_{\theta}^{\Phi}$.
Assumption 2.4.
$\mathcal{X}\subset\mathbb{R}^{n}$ and $\mathcal{Y}\subset\mathbb{R}^{d}$ are a convex and compact sets with diameters $D_{\mathcal{X}}$ and $D_{\mathcal{Y}}$, respectively, i.e., $D_{\mathcal{X}}\triangleq sup_{x,\overline{x}\in\mathcal{X}}\left\|x-\overline{x}\right\|_{\mathcal{X}}$ and $D_{\mathcal{Y}}\triangleq sup_{y,\overline{y}\in\mathcal{Y}}\left\|y-\overline{y}\right\|_{\mathcal{Y}}$.
In this paper, to handle the primal constraint set $\mathcal{X}$ we use two different oracles, namely a linear minimization oracle (LMO) to find $s=\operatorname*{\arg\!\min}_{x\in\mathcal{X}}\left\langle x,\bar{x}\right\rangle$ for any given $\bar{x}\in\mathbb{R}^{n}$ and a projection oracle (PO) to find $s=\operatorname*{\arg\!\min}_{x\in\mathcal{X}}\left\|x-\bar{x}\right\|$ for any given $\bar{x}\in\mathbb{R}^{n}$. Consequently, to measure the quality of solutions generated by the proposed algorithm we define two different gap functions. The first gap function will utilize LMO for the primal part and the generalized gradient norm for the dual part while the second one will use the generalized gradient norm for both the primal and dual parts.
Definition 2.1.
The stationary gap function $\mathcal{G}_{\mathcal{X}}:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ for the primal part of Problem (1) is defined as $\mathcal{G}_{\mathcal{X}}(x,y)\triangleq\sup_{s\in\mathcal{X}}\langle\nabla_{x}\mathcal{L}(x,y),x-s\rangle$. For the dual part of Problem (1), the stationary gap function $\mathcal{G}_{\mathcal{Y}}:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ is defined as $\mathcal{G}_{\mathcal{Y}}(x,y)\triangleq\frac{1}{\sigma}\left\|y-\mathcal{P}_{\mathcal{Y}}(y+\sigma\nabla_{y}\mathcal{L}(x,y))\right\|$. Moreover, we define $\mathcal{G}_{\mathcal{Z}}(x,y)\triangleq\mathcal{G}_{\mathcal{X}}(x,y)+\mathcal{G}_{\mathcal{Y}}(x,y)$.
Definition 2.2.
The stationary gap function $\mathcal{G}_{\mathcal{X}}:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ for the primal part of Problem (1) is defined as $\mathcal{G}_{\mathcal{X}}(x,y)\triangleq\frac{1}{\tau}\left\|s-x\right\|$. The dual gap function in Algorithm 2 is defined similarly to the dual function of Algorithm 1. Additionally, we define $\mathcal{G}_{\mathcal{Z}}(x,y)\triangleq\mathcal{G}_{\mathcal{X}}(x,y)+\mathcal{G}_{\mathcal{Y}}(x,y)$ where $\mathcal{G}_{\mathcal{Y}}$ is introduced in Definition 2.1.
Definition 2.3.
For a given gap function $\mathcal{G}_{\mathcal{Z}}:\mathcal{Z}\rightarrow\mathbb{R}$, where $\mathcal{Z}\triangleq\mathcal{X}\times\mathcal{Y}$, we define a point $(x,y)\in\mathcal{Z}$ as an $\epsilon$-stationary solution for Problem (1) if $\mathcal{G}_{\mathcal{Z}}(x,y)\leq\epsilon$.
Remark 2.2.
With a slight abuse of notation, we utilized $\mathcal{G}_{\mathcal{X}}$ for both the Linear Minimization Oracle (LMO) and the Projection Oracle (PO) with respect to the constraint set $\mathcal{X}$. In fact, both $\epsilon$-solutions imply an $\epsilon$-game stationary solution. Specifically, if $\mathcal{G}_{\mathcal{Z}}(x,y)\leq\epsilon$ holds for some $(x,y)\in\mathcal{Z}$, then
$0\in\partial\mathds{1}_{\mathcal{X}}(x)+\nabla_{x}\mathcal{L}(x,y)+u,\quad\text{s.t.}\quad u\in\mathcal{B}_{\mathbb{R}^{n}}(0,\epsilon),$
$0\in\partial\mathds{1}_{\mathcal{Y}}(y)-\nabla_{y}\mathcal{L}(x,y)+v,\quad\text{s.t.}\quad v\in\mathcal{B}_{\mathbb{R}^{d}}(0,\epsilon).$
where we define $\mathcal{B}_{\mathcal{U}}(u,r)\triangleq\big{\{}u\in\mathcal{U}\big{|}\|u\|\leq r\big{\}}$ for a given vector space $\mathcal{U}$.
2.2 Problem Properties
To solve Problem (1), the initial step involves constructing two implicit gradients for the SP objective function: one corresponding to the minimization variable $x$ and the other for the maximization variable $y$. More specifically,
$\nabla_{x}\mathcal{L}(x,y)=\nabla_{x}\Phi(x,\theta^{*}(x),y)+{\bf J}\theta^{*}(x)\nabla_{\theta}\Phi(x,\theta^{*}(x),y)$ (6a)
$\nabla_{y}\mathcal{L}(x,y)=\nabla_{y}\Phi(x,\theta^{*}(x),y)$ (6b)
where ${\bf J}\theta^{*}(x)\in\mathbb{R}^{m\times n}$ is the Jacobian of $\theta^{*}(x)$. One can characterize the Jacobian matrix based on the implicit function theorem for the lower-level problem by writing its first-order optimality condition, namely $\nabla_{\theta}g(x,\theta^{*}(x))=0$ for any $x\in\mathbb{R}^{n}$. By taking the derivative with respect to $x$, one can obtain ${\bf J}\theta^{*}(x)=-\nabla_{\theta x}^{2}g(x,\theta^{*}(x))[\nabla_{\theta\theta}^{2}g(x,\theta^{*}(x))]^{-1}$. Substituting this relation in (6a), we have
$\nabla_{x}\mathcal{L}(x,y)=\nabla_{x}\Phi(x,\theta^{*}(x),y)-\nabla_{\theta x}^{2}g(x,\theta^{*}(x))[\nabla_{\theta\theta}^{2}g(x,\theta^{*}(x))]^{-1}\nabla_{\theta}\Phi(x,\theta^{*}(x),y)$ (7)
which can be rewritten as follows
$\nabla_{x}\mathcal{L}(x,y)=\nabla_{x}\Phi(x,\theta^{*}(x),y)-\nabla_{\theta x}^{2}g(x,\theta^{*}(x))v(x,y),$ (8a)
$\text{where}\quad v(x,y)=[\nabla_{\theta\theta}^{2}g(x,\theta^{*}(x))]^{-1}\nabla_{\theta}\Phi(x,\theta^{*}(x),y).$ (8b)
The above formulation presents several challenges inherent to bilevel optimization problems, notably finding $\theta^{*}(x)$ and the inversion of the Hessian matrix. Furthermore, our SP structure adds another layer of complexity compared to typical bilevel optimization setups. This complexity stems from the interplay, both explicit and implicit, between the maximization variable $y$, the primal variable $x$, and the lower-level optimal solution $\theta^{*}(x)$, as represented in equation (6b).
Based on the considered setting and assumptions, we first establish several important properties of the problem. In particular, we first establish the Lipschitz continuity of map $v$ in both primal and dual variables. Consequently, we can show the Lipschitz continuity of the lower-level optimal solution and partial derivatives of the implicit objective function $\mathcal{L}$. These results serve as the cornerstone for the analysis of our proposed methods in the following section.
Lemma 2.1.
Suppose Assumptions 2.1, 2.2 hold. Then for any $x,\overline{x}\in\mathcal{X}$ and $y,\overline{y}\in\mathcal{Y}$, $\left\|v(x,y)-v(\overline{x},\overline{y})\right\|\leq\mathbf{C}_{v1}\left\|x-\overline{x}\right\|+\mathbf{C}_{v2}\left\|y-\overline{y}\right\|$ where $\mathbf{C}_{v1}=\frac{L_{\theta x}^{\Phi}+L_{\theta\theta}^{\Phi}\mathbf{L}_{\theta}}{\mu_{g}}+\frac{C_{\theta}^{\Phi}L_{\theta\theta}^{g}}{\mu^{2}_{g}}(1+\mathbf{L}_{\theta})$ and $\mathbf{C}_{v2}=\frac{1}{\mu_{g}}L_{\theta y}^{\Phi}$.
Lemma 2.2.
Suppose Assumptions 2.1, 2.2 hold. Then for any $x,\overline{x}\in\mathcal{X}$ and $y,\overline{y}\in\mathcal{Y}$, the following results hold.
  1. (I)
    There exists $\mathbf{L}_{\theta}\geq 0$ such that $\left\|\theta^{*}(x)-\theta^{*}(\overline{x})\right\|\leq\mathbf{L}_{\theta}\left\|x-\overline{x}\right\|$ where $\mathbf{L}_{\theta}=\frac{C_{\theta x}^{g}}{\mu_{g}}$.
  2. (II)
    There exists $\mathbf{L}_{\mathcal{L}_{1}},\mathbf{L}_{\mathcal{L}_{2}}\geq 0$ such that $\left\|\nabla_{x}\mathcal{L}(x,y)-\nabla_{x}\mathcal{L}(\overline{x},\overline{y})\right\|\leq\mathbf{L}_{\mathcal{L}_{1}}\left\|x-\overline{x}\right\|+\mathbf{L}_{\mathcal{L}_{2}}\left\|y-\overline{y}\right\|$ where $\mathbf{L}_{\mathcal{L}_{1}}=(L_{xx}^{\Phi}+L_{x\theta}^{\Phi}\mathbf{L}_{\theta}+C_{\theta x}^{g}\mathbf{C}_{v1}+\frac{C_{\theta}^{\Phi}}{\mu_{g}}L_{\theta x}^{g}(1+\mathbf{L}_{\theta})),\mathbf{L}_{\mathcal{L}_{2}}=(L_{xy}^{\Phi}+C_{\theta x}^{g}\mathbf{C}_{v2})$.
  3. (III)
    $\|\nabla_{y}\mathcal{L}(x,y)-\nabla_{y}\mathcal{L}(\overline{x},\overline{y})\|\leq(L_{yx}^{\Phi}+L_{y\theta}^{\Phi}\mathbf{L}_{\theta})\left\|x-\overline{x}\right\|+L_{yy}^{\Phi}\left\|y-\overline{y}\right\|$ where $\mathbf{L}_{\theta}=\frac{C_{\theta x}^{g}}{\mu_{g}}$.
3 Proposed Methods
In this section, we present our proposed algorithms based on an inexact bilevel regularized primal-dual approach for solving Problem (1). We consider two main approaches to deal with minimization constraint $\mathcal{X}$: the first one involves using a Linear Minimization Oracle (LMO), while the second in the approach we use a Projection Oracle (PO) to update the primal step of the upper-level objective function. To update the maximization variable we consider a gradient ascent-based approach. Moreover, in the lower-level problem, we incorporate one step of gradient-type update to estimate the lower-level optimal solution. As a result, we propose both a one-sided projection-free and a fully projected gradient-based algorithm. Next, we discuss the detailed update of our proposed methods.
Primal Update: one common approach to avoid using Hessian inverse in (8b) is to solve the corresponding parametric quadratic programming problem as follows
$v(x,y)=\operatorname*{\arg\!\min}_{v}\frac{1}{2}v^{\top}\nabla_{\theta\theta}^{2}g(x,\theta^{*}(x))v-\nabla_{\theta}\Phi(x,\theta^{*}(x),y)^{\top}v.$ (9)
The solution to the above quadratic programming problem can be approximated using one step of the gradient descent technique [30, 35], as follows:
$v(x_{k},y_{k})=v(x_{k},y_{k})-\eta_{k}\big{(}\nabla_{\theta\theta}^{2}g(x_{k},\theta^{*}(x_{k}))v(x_{k},y_{k})-\nabla_{\theta}\Phi(x_{k},\theta^{*}(x_{k}),y_{k})\big{)}.$
for some step-size $\eta_{k}\geq 0$. In order to implement an iterative method for solving the SPB problem using first-order information at each iteration $k\geq 0$, the exact evaluation of $\theta^{*}(x_{k})$ is typically not feasible. One approach is to substitute $\theta^{*}(x_{k})$ with its estimated solution $\theta_{k}$, which again can be obtained by applying one step of the gradient descent method with respect to the lower-level objective function to trace the optimal trajectory of the lower-level problem. Subsequently, $v(x_{k},y_{k})$ can be approximated by an iterate $w_{k+1}$ obtained by taking one step of gradient descent as
$w_{k+1}\leftarrow w_{k}-\eta_{k}\big{(}\nabla_{\theta\theta}^{2}g(x_{k},\theta_{k})w_{k}-\nabla_{\theta}\Phi(x_{k},\theta_{k},y_{k})\big{)}.$
Next, the directions estimating (8a) and (6b) can be obtained as
$G^{x}_{k}=\nabla_{x}\Phi(x,\theta_{k},y)-\nabla_{\theta x}^{2}g(x,\theta_{k})w_{k+1},$
$G^{y}_{k}=\nabla_{y}\Phi(x,\theta_{k},y).$
Then, based on the oracle available for the constraint set $\mathcal{X}$, we either employ the Frank-Wolfe (FW)-type update [14] or projected gradient (PG) update based on the Krasnoselskii-Mann method [48]. In particular, we compute the next iteration $x_{k+1}$ using FW and PG, respectively, as follows:
$FW:s_{k}\leftarrow\operatorname*{\arg\!\min}_{x\in\mathcal{X}}\langle G_{k}^{x},x\rangle,\quad x_{k+1}\leftarrow\gamma_{k}s_{k}+(1-\gamma_{k})x_{k},$
$PG:s_{k}\leftarrow\mathcal{P}_{\mathcal{X}}(x_{k}-\tau_{k}G_{k}^{x}),\quad x_{k+1}\leftarrow\gamma_{k}s_{k}+(1-\gamma_{k})x_{k}.$
for some step-size $\gamma_{k}\in[0,1]$. Given the update $x_{k+1}$, by implementing a gradient descent step with respect to the lower-level objective function $g(x_{k+1},\cdot)$ using a step-size $\alpha>0$, we can generate an iterative sequence $\theta_{k+1}$ to approximate the lower-level optimal solution $\theta^{*}(x_{k+1})$ as
$\theta_{k+1}\leftarrow\theta_{k}-\alpha\nabla_{\theta}g(x_{k+1},\theta_{k}).$
Dual Update: In the context of nonconvex-concave (NC-C) SP problem, many approaches have been introduced recently to mitigate the challenge of nonsmoothness and nonconvexity of $f(x)\triangleq\max_{y\in\mathcal{Y}}\mathcal{L}(x,y)$. One approach is dual smoothing in which we add a regularization term $-\frac{\mu}{2}\|y-y_{0}\|^{2}$ for some $y_{0}\in\mathcal{Y}$, to ensure that $f$ is smooth [58]. Therefore, we can write the regularized upper-level objective function for Problem (1) as follows
$\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\mathcal{L}_{\mu}(x,y)\triangleq\mathcal{L}(x,y)-\frac{\mu}{2}\|y-y_{0}\|^{2}.$
Moreover, suppose $f_{\mu}:\mathcal{X}\rightarrow\mathbb{R}$ be a function such that $f_{\mu}(x)\triangleq\max_{y\in\mathcal{Y}}\mathcal{L}(x,y)-\frac{\mu}{2}\left\|y-y_{0}\right\|^{2}$, we also could define $y^{*}_{\mu}(x)\triangleq\operatorname*{\arg\!\min}_{y\in\mathcal{Y}}\mathcal{L}(x,y)-\frac{\mu}{2}\left\|y-y_{0}\right\|^{2}$. Here, by controlling $\mu$, we can find $y^{*}_{\mu}(x)$ close to $y^{*}(x)$ (i.e., $y^{*}_{\mu}(x)\approx y^{*}(x)$). Consequently, we can update the dual variable by performing a step of projected gradient ascent, as follows:
$y_{k+1}\leftarrow\mathcal{P}_{\mathcal{Y}}\Big{(}y_{k}+\sigma_{k}\big{(}G_{k}^{y}-\mu(y_{k}-y_{0})\big{)}\Big{)}$
Our proposed inexact bilevel regularized primal-dual algorithms, one-sided projection-free (i-BRPD:OPF) and fully projected (i-BRPD:FP), are outlined in Algorithms 1 and 2, respectively.
Algorithm 1 Inexact Bilevel Regularized Primal-dual: One-sided Projection-free (i-BRPD:OPF) Method
1:  Input: $x_{0}\in\mathcal{X}$, $y_{0}\in\mathcal{Y}$, $\theta_{0}\in\mathbb{R}^{m}$, $\alpha,\mu>0$, $\{\gamma_{k},\sigma_{k},\eta_{k}\}_{k}\subseteq\mathbb{R}_{+}$
2:  Initialization: $\mathbf{w}^{0}\leftarrow\boldsymbol{\theta}^{0}$
3:  for $k=0,\dots,K-1$ do
4:     $w_{k+1}\leftarrow w_{k}-\eta_{k}(\nabla_{\theta\theta}^{2}g(x_{k},\theta_{k})w_{k}-\nabla_{\theta}\Phi(x_{k},\theta_{k},y_{k}))$
5:     $G_{k}^{x}\leftarrow\nabla_{x}\Phi(x_{k},\theta_{k},y_{k})-\nabla_{\theta x}^{2}g(x_{k},\theta_{k})w_{k+1}$
6:     $G_{k}^{y}\leftarrow\nabla_{y}\Phi(x_{k},\theta_{k},y_{k})$
7:     $s_{k}\leftarrow\operatorname*{\arg\!\min}_{x\in\mathcal{X}}\langle G_{k}^{x},x\rangle$
8:     $x_{k+1}\leftarrow\gamma_{k}s_{k}+(1-\gamma_{k})x_{k}$
9:     $\theta_{k+1}\leftarrow\theta_{k}-\alpha\nabla_{\theta}g(x_{k+1},\theta_{k})$
10:     $y_{k+1}\leftarrow\mathcal{P}_{\mathcal{Y}}\Big{(}y_{k}+\sigma_{k}\big{(}G_{k}^{y}-\mu(y_{k}-y_{0})\big{)}\Big{)}$
11:  end for
Algorithm 2 Inexact Bilevel Regularized Primal-dual: Fully Projected (i-BRPD:FP) Method
1:  Input: $x_{0}\in\mathcal{X}$, $y_{0}\in\mathcal{Y}$, $\theta_{0}\in\mathbb{R}^{m}$, $\alpha,\mu>0$, $\{\gamma_{k},\sigma_{k},\tau_{k},\eta_{k}\}_{k}\subseteq\mathbb{R}_{+}$
2:  Initialization: $\mathbf{w}^{0}\leftarrow\boldsymbol{\theta}^{0}$
3:  for $k=0,\dots,K-1$ do
4:     $w_{k+1}\leftarrow w_{k}-\eta_{k}(\nabla_{\theta\theta}^{2}g(x_{k},\theta_{k})w_{k}-\nabla_{\theta}\Phi(x_{k},\theta_{k},y_{k}))$
5:     $G_{k}^{x}\leftarrow\nabla_{x}\Phi(x_{k},\theta_{k},y_{k})-\nabla_{\theta x}^{2}g(x_{k},\theta_{k})w_{k+1}$
6:     $G_{k}^{y}\leftarrow\nabla_{y}\Phi(x_{k},\theta_{k},y_{k})$
7:     $s_{k}\leftarrow\mathcal{P}_{\mathcal{X}}(x_{k}-\tau_{k}G_{k}^{x})$
8:     $x_{k+1}\leftarrow\gamma_{k}s_{k}+(1-\gamma_{k})x_{k}$
9:     $\theta_{k+1}\leftarrow\theta_{k}-\alpha\nabla_{\theta}g(x_{k+1},\theta_{k})$
10:     $y_{k+1}\leftarrow\mathcal{P}_{\mathcal{Y}}\Big{(}y_{k}+\sigma_{k}\big{(}G_{k}^{y}-\mu(y_{k}-y_{0})\big{)}\Big{)}$
11:  end for
It is noteworthy to mention that our proposed algorithms are designed as single-loop algorithms, making them easy to implement. Moreover, at each iteration, we need at most two Hessian vector products.
4 Analysis of the Proposed Methods
In this section, we first delve into the analysis of the lower-level approximation in Subsection 4.1. Moving on to Subsection 4.2, we establish a bound on the error between $y_{k}$ and $y_{\mu}^{*}(x_{k},y_{k})$ to scrutinize the trajectory of the optimal solution in the maximization component. Bringing together these subsections, we elucidate the analysis of gradient estimation in Subsection 4.3.
4.1 Analysis of the Lower-level Approximation
The first step is to analyze the lower-level to see how $\theta_{k}$ is tracking $\theta^{*}(x_{k})$, which is crucial for understanding the convergence behavior of Algorithm 1. This analysis serves to establish bounds on the error between $\theta_{k}$ and $\theta^{*}(x_{k})$, which is essential for assessing the convergence rate of the algorithm.
Lemma 4.1.
Suppose Assumption 2.2 holds. Let $\{(x_{k},\theta_{k})\}_{k\geq 0}$ be the sequence generated by Algorithm 1, such that $\alpha=\frac{2}{\mu_{g}+L_{g}}$. Then, for any $k\geq 0$
$\left\|\theta_{k}-\theta^{*}(x_{k})\right\|\leq\beta^{k}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\mathbf{L}_{\theta}\sum_{i=0}^{k-1}\beta^{k-i}\left\|x_{i}-x_{i+1}\right\|$ (10)
where $\beta\triangleq(L_{g}-\mu_{g})/(L_{g}+\mu_{g})$.
Proof.
We begin the proof by characterizing the one-step progress of the lower-level iterate sequence $\{\boldsymbol{\theta}_{k}\}_{k}$. Indeed, at iteration $k$, we aim to approximate $\boldsymbol{\theta}^{*}(\mathbf{x}_{k+1})=\operatorname*{\arg\!\min}_{\boldsymbol{\theta}}g(\mathbf{x}_{k+1},\boldsymbol{\theta})$. According to the update of $\boldsymbol{\theta}_{k+1}$, we observe that
$\left\|\theta_{k+1}-\theta^{*}(x_{k+1})\right\|^{2}$ $=\left\|\theta_{k}-\theta^{*}(x_{k+1})-\alpha\nabla_{\theta}g(x_{k+1},\theta_{k})\right\|^{2}$
$=\left\|\theta_{k}-\theta^{*}(x_{k+1})\right\|^{2}-2\alpha\langle\nabla_{\theta}g(x_{k+1},\theta_{k}),\theta_{k}-\theta^{*}(x_{k+1})\rangle$
$\quad+\alpha^{2}\left\|\nabla_{\theta}g(x_{k+1},\theta_{k})\right\|^{2}.$ (11)
Moreover, from Assumption 2.2 we have that
$\langle\nabla_{\theta}g(x_{k+1},\theta_{k}),\theta_{k}-\theta^{*}(x_{k+1})\rangle\geq\frac{\mu_{g}L_{g}}{\mu_{g}+L_{g}}\left\|\theta_{k}-\theta^{*}(x_{k+1})\right\|^{2}+\frac{1}{\mu_{g}+L_{g}}\left\|\nabla_{\theta}g(x_{k+1},\theta_{k})\right\|^{2}.$ (12)
The inequality in (4.1) together with (12) imply that
$\left\|\theta_{k+1}-\theta^{*}(x_{k+1})\right\|^{2}$ $\leq\left\|\theta_{k}-\theta^{*}(x_{k+1})\right\|^{2}-\frac{2\alpha\mu_{g}L_{g}}{\mu_{g}+L_{g}}\left\|\theta_{k}-\theta^{*}(x_{k+1})\right\|^{2}$
$\quad+(\alpha^{2}-\frac{2\alpha}{\mu_{g}+L_{g}})\left\|\nabla_{\theta}g(x_{k+1},\theta_{k}\right\|^{2}.$ (13)
Setting the step-size $\alpha=\frac{2}{\mu_{g}+L_{g}}$ in (4.1) leads to
$\left\|\theta_{k+1}-\theta^{*}(x_{k+1})\right\|^{2}\leq(\frac{\mu_{g}-L_{g}}{\mu_{g}+L_{g}})^{2}\left\|\theta_{k}-\theta^{*}(x_{k+1})\right\|^{2}$ (14)
Next, recall that $\beta=(L_{g}-\mu_{g})/(L_{g}+\mu_{g})$. Using the triangle inequality and Part (I) of Lemma 2.2 we conclude that
$\left\|\theta_{k+1}-\theta^{*}(x_{k+1})\right\|$ $\leq\beta\left\|\theta_{k}-\theta^{*}(x_{k+1})\right\|$
$\leq\beta\big{[}\left\|\theta_{k}-\theta^{*}(x_{k})\right\|+\left\|\theta^{*}(x_{k})-\theta^{*}(x_{k+1})\right\|\big{]}$
$\leq\beta\big{[}\left\|\theta_{k}-\theta^{*}(x_{k})\right\|+\mathbf{L}_{\theta}\left\|x_{k}-x_{k+1}\right\|\big{]}.$ (15)
Moreover, from the update of $\mathbf{x}_{k+1}$ in Algorithm 1 and boundedness of $\mathcal{X}$ we have that $\left\|\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\|\leq\gamma_{k}D_{\mathcal{X}}$. Therefore, using this inequality within (4.1) leads to
$\left\|\boldsymbol{\theta}_{k+1}-\boldsymbol{\theta}^{*}(\mathbf{x}_{k+1})\right\|\leq\beta\left\|\boldsymbol{\theta}_{k}-\boldsymbol{\theta}^{*}(\mathbf{x}_{k})\right\|+\beta\gamma_{k}D_{\mathcal{X}}\mathbf{L}_{\theta}.$
Finally, the desired result can be deduced from the above inequality recursively.
In the next step of analyzing the lower-level approximation, we need to understand how well $w_{k+1}$ tracks $v(x_{k},y_{k})$. As explained in Section 3, $v(x_{k},y_{k})$ can be approximated by an iterate $w_{k+1}$ obtained by taking one step of gradient descent. The analysis of the bound on the error between $w_{k+1}$ and $v(x_{k},y_{k})$ is crucial for understanding the behavior and convergence properties of the proposed algorithms.
Lemma 4.2.
Let $\{(x_{k},w_{k},y_{k})\}_{k\geq 0}$ be the sequence generated by the Algorithm 1 or 2 and suppose Assumptions 2.1, 2.2 hold and $\eta\leq 2/(L_{g}+\mu_{g})$. Then, for any $k\geq 0$,
$\|w_{k+1}-v(x_{k},y_{k})\|$ $\leq\rho^{k+1}\|w_{0}-v(x_{0},y_{0})\|+\gamma\mathbf{C}_{v1}\sum_{i=0}^{k-1}\rho^{k-i}\|s_{i}-x_{i}\|+\mathbf{C}_{v2}\sum_{i=0}^{k-1}\rho^{k-i}\|y_{i+1}-y_{i}\|$
$\quad+\tilde{C}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|k\rho^{k}+\gamma\tilde{C}\mathbf{L}_{\theta}\sum_{i=0}^{k}\sum_{j=0}^{i-1}\rho^{k-i}\beta^{i-j}\left\|s_{j}-x_{j}\right\|$ (16)
where $\tilde{C}=\eta(L_{\theta\theta}^{g}\frac{C_{\theta}^{\Phi}}{\mu_{g}}+L_{\theta\theta}^{\Phi})$ and $\rho\triangleq(1-\eta\mu_{g})$.
Proof.
By verifying that $v(x_{k},y_{k})=v(x_{k},y_{k})-\eta\big{(}\nabla_{\theta\theta}^{2}g(x_{k},\theta^{*}(x_{k}))v(x_{k},y_{k})-\nabla_{\theta}\Phi(x_{k},\theta^{*}(x_{k}),y_{k})\big{)}$, and by defining $w_{k+1}=w_{k}-\eta\big{(}\nabla_{\theta\theta}^{2}g(x_{k},\theta_{k})w_{k}-\nabla_{\theta}\Phi(x_{k},\theta_{k},y_{k})\big{)}$ we can write
$\|w_{k+1}-v(x_{k},y_{k})\|$ $=\|\Big{(}w_{k}-\eta\big{(}\nabla_{\theta\theta}^{2}g(x_{k},\theta_{k})w_{k}-\nabla_{\theta}\Phi(x_{k},\theta_{k},y_{k})\big{)}\Big{)}$
$\quad-\Big{(}v(x_{k},y_{k})-\eta\big{(}\nabla_{\theta\theta}^{2}g(x_{k},\theta^{*}(x_{k}))v(x_{k},y_{k})-\nabla_{\theta}\Phi(x_{k},\theta^{*}(x_{k}),y_{k})\big{)}\Big{)}\|$
$=\|\Big{(}I-\eta\nabla_{\theta\theta}^{2}g(x_{k},\theta_{k})\Big{)}(w_{k}-v(x_{k},y_{k}))-\eta\Big{(}\nabla_{\theta\theta}^{2}g(x_{k},\theta_{k})$
$\quad-\nabla_{\theta\theta}^{2}g(x_{k},\theta^{*}(x_{k}))\Big{)}v(x_{k},y_{k})+\eta\Big{(}\nabla_{\theta}\Phi(x_{k},\theta^{*}(x_{k}),y_{k})-\nabla_{\theta}\Phi(x_{k},\theta_{k},y_{k})\Big{)}\|,$ (17)
where the last equality is obtained by adding and subtracting the term $(I-\eta\nabla_{\theta\theta}^{2}g(x_{k},\theta_{k}))v(x_{k},y_{k})$. Next, using Assumptions 2.1 and 2.2 along with the application of the triangle inequality we obtain
$\|w_{k+1}-v(x_{k},y_{k})\|$ $\leq(1-\eta\mu_{g})\|w_{k}-v(x_{k},y_{k})\|+\eta L_{\theta\theta}^{g}\|\theta_{k}-\theta^{*}(x_{k})\|\|v(x_{k},y_{k})\|$
$\quad+\eta L_{\theta\theta}^{\Phi}\|\theta_{k}-\theta^{*}(x_{k})\|.$ (18)
Note that $\|v(x,y)\|=\|[\nabla_{\theta\theta}^{2}g(x,\theta^{*}(x))]^{-1}\nabla_{\theta}\Phi(x,\theta^{*}(x),y)\|\leq\frac{C_{\theta}^{\Phi}}{\mu_{g}}$. Now, by adding and subtracting $v(x_{k-1},y_{k-1})$ to the term $\|w_{k}-v(x_{k},y_{k})\|$ followed by the triangle inequality application we can conclude that
$\|w_{k+1}-v(x_{k},y_{k})\|$ $\leq(1-\eta\mu_{g})\|w_{k}-v(x_{k-1},y_{k-1})\|+(1-\eta\mu_{g})\|v(x_{k-1},y_{k-1})-v(x_{k},y_{k})\|$
$\quad+\eta(L_{\theta\theta}^{g}\frac{C_{\theta}^{\Phi}}{\mu_{g}}+L_{\theta\theta}^{\Phi})\|\theta_{k}-\theta^{*}(x_{k})\|.$ (19)
Therefore, using the result of Lemma 2.1 and 4.1 we can show
$\|w_{k+1}-v(x_{k},y_{k})\|$ $\leq\rho\|w_{k}-v(x_{k-1},y_{k-1})\|+\rho(\mathbf{C}_{v1}\left\|x_{i}-x_{i-1}\right\|+\mathbf{C}_{v2}\left\|y_{i}-y_{i-1}\right\|)$
$\quad+\eta(L_{\theta\theta}^{g}\frac{C_{\theta}^{\Phi}}{\mu_{g}}+L_{\theta\theta}^{\Phi})(\beta^{i}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\mathbf{L}_{\theta}\sum_{j=0}^{i-1}\beta^{i-j}\left\|x_{j}-x_{j+1}\right\|).$ (20)
Consequently, by continuing the above inequality recursively from 00 to $k$, we obtain
$\|w_{k+1}-v(x_{k},y_{k})\|$ $\leq\rho^{k+1}\|w_{0}-v(x_{0},y_{0})\|+\sum_{i=0}^{k}\rho^{k-i+1}(\mathbf{C}_{v1}\left\|x_{i}-x_{i-1}\right\|+\mathbf{C}_{v2}\left\|y_{i}-y_{i-1}\right\|)$
$\quad+\sum_{i=0}^{k}\rho^{k-i}\eta(L_{\theta\theta}^{g}\frac{C_{\theta}^{\Phi}}{\mu_{g}}+L_{\theta\theta}^{\Phi})(\beta^{i}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\mathbf{L}_{\theta}\sum_{j=0}^{i-1}\beta^{i-j}\left\|x_{j}-x_{j+1}\right\|)$
$\leq\rho^{k+1}\|w_{0}-v(x_{0},y_{0})\|+\mathbf{C}_{v1}\sum_{i=0}^{k-1}\rho^{k-i}\|x_{i+1}-x_{i}\|+\mathbf{C}_{v2}\sum_{i=0}^{k-1}\rho^{k-i}\|y_{i+1}-y_{i}\|$
$\quad+\tilde{C}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|\sum_{i=0}^{k}\rho^{k-i}\beta^{i}+\tilde{C}\mathbf{L}_{\theta}\sum_{i=0}^{k}\rho^{k-i}\sum_{j=0}^{i-1}\beta^{i-j}\left\|x_{j}-x_{j+1}\right\|$
$\leq\rho^{k+1}\|w_{0}-v(x_{0},y_{0})\|+\mathbf{C}_{v1}\sum_{i=0}^{k-1}\rho^{k-i}\|x_{i+1}-x_{i}\|+\mathbf{C}_{v2}\sum_{i=0}^{k-1}\rho^{k-i}\|y_{i+1}-y_{i}\|$
$\quad+\tilde{C}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|k\rho^{k}+\tilde{C}\mathbf{L}_{\theta}\sum_{i=0}^{k}\sum_{j=0}^{i-1}\rho^{k-i}\beta^{i-j}\left\|x_{j}-x_{j+1}\right\|$
where $\tilde{C}=\eta(L_{\theta\theta}^{g}\frac{C_{\theta}^{\Phi}}{\mu_{g}}+L_{\theta\theta}^{\Phi})$ and in the last inequality we used the fact that $\beta\leq\rho$. The result follows by rearranging the terms on the RHS and $x_{k+1}-x_{k}=\gamma(s_{k}-x_{k})$ for any $k\geq 0$.
4.2 Analysis of Tracking the Optimal Solution Trajectory of the Maximization Component
In the first step of examining the trajectory of the optimal solution in the maximization component, we need to discuss two important properties outlined in Lemmas 4.3 and 4.4, which have already been proven in [4] within the context of the SP problem. Here, we present the lemmas and their proofs for completeness.
Lemma 4.3.
The solution map $y^{*}_{\mu}:\mathcal{X}\rightarrow\mathcal{Y}$ is Lipschitz continuous. Suppose Assumption 2.1, 2.2 hold and $\theta^{*}(x)$ is Lipschitz continuous map with constant $\mathbf{L}_{\theta}$. In particular, for any $x,\overline{x}\in\mathcal{X}$
$\left\|y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x)\right\|\leq\frac{L_{y\mu}}{\mu}\left\|x-\overline{x}\right\|$ (21)
where $L_{y\mu}=L_{yx}^{\Phi}+\mathbf{L}_{\theta}L_{y\theta}^{\Phi}$.
Proof.
Since $\mathcal{L}_{\mu}(x,\cdot)$ is strongly concave for any $x\in\mathcal{X}$, we have that
$(y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x))^{T}\big{[}\nabla_{y}\mathcal{L}_{\mu}(x,y^{*}_{\mu}(\overline{x}))-\nabla_{y}\mathcal{L}_{\mu}(x,y^{*}_{\mu}(x))\big{]}+\mu\left\|y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x)\right\|^{2}\leq 0.$ (22)
Moreover, the optimality of $y^{*}_{\mu}(\overline{x})$ and $y^{*}_{\mu}(x)$ given that $\mathcal{L}_{\mu}(x,y)=\mathcal{L}(x,y)-\frac{\mu}{2}\left\|y-y_{0}\right\|^{2}$ implies that for any $y\in\mathcal{Y}$,
$(y-y^{*}_{\mu}(\overline{x}))^{T}\nabla_{y}\mathcal{L}_{\mu}(\overline{x},y^{*}_{\mu}(\overline{x}))\leq 0,$ (23)
$(y-y^{*}_{\mu}(x))^{T}\nabla_{y}\mathcal{L}_{\mu}(x,y^{*}_{\mu}(x))\leq 0.$ (24)
Let $y=y^{*}_{\mu}(x)$ in (23) and $y=y^{*}_{\mu}(\overline{x})$ in (24) and summing up two inequalities, we obtain
$(y^{*}_{\mu}(x)-y^{*}_{\mu}(\overline{x}))^{T}\Big{(}\nabla_{y}\mathcal{L}_{\mu}(\overline{x},y^{*}_{\mu}(\overline{x}))-\nabla_{y}\mathcal{L}_{\mu}(x,y^{*}_{\mu}(x))\Big{)}\leq 0.$ (25)
By combining (25) and (22) we have
$\mu\left\|y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x)\right\|^{2}\leq(y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x))^{T}\Big{(}\nabla_{y}\mathcal{L}_{\mu}(\overline{x},y^{*}_{\mu}(\overline{x}))-\nabla_{y}\mathcal{L}_{\mu}(x,y^{*}_{\mu}(\overline{x}))\Big{)}$ (26)
We know from (6b) that $\nabla_{y}\mathcal{L}(x,y)=\nabla_{y}\Phi(x,\theta^{*}(x),y)$ and by defining $\Phi_{\mu}\triangleq\Phi-\frac{\mu}{2}\left\|y-y_{0}\right\|^{2}$ we have
$\mu\left\|y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x)\right\|^{2}\leq(y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x))^{T}\Big{(}\nabla_{y}\Phi_{\mu}(\overline{x},\theta^{*}(\overline{x}),y^{*}_{\mu}(\overline{x}))-\nabla_{y}\Phi_{\mu}(x,\theta^{*}(x),y^{*}_{\mu}(\overline{x}))\Big{)}$ (27)
Using Assumption 2.1-(3) we can get
$\mu\left\|y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x)\right\|^{2}\leq\left\|y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x)\right\|(L_{yx}^{\Phi}\left\|x-\overline{x}\right\|+L_{y\theta}^{\Phi}\left\|\theta^{*}(x)-\theta^{*}(\overline{x})\right\|)$ (28)
$\left\|y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x)\right\|\leq\frac{L_{yx}^{\Phi}}{\mu}\left\|x-\overline{x}\right\|+\frac{L_{y\theta}^{\Phi}}{\mu}\left\|\theta^{*}(x)-\theta^{*}(\overline{x})\right\|$ (29)
Considering Lemma 2.2-(I) we can show that
$\left\|y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x)\right\|\leq(\frac{L_{yx}^{\Phi}+\mathbf{L}_{\theta}L_{y\theta}^{\Phi}}{\mu})\left\|x-\overline{x}\right\|$
Lemma 4.4.
Suppose Assumption 2.1, 2.2 hold. The function $f_{\mu}(\cdot)$ is differentiable on an open set containing $\mathcal{X}$ and $\nabla f_{\mu}(x)=\nabla_{x}\mathcal{L}(x,y^{*}_{\mu}(x))$ where $y^{*}_{\mu}(x)\triangleq\operatorname*{\arg\!\min}_{y\in\mathcal{Y}}\mathcal{L}_{\mu}(x,y)$. Moreover, $f_{\mu}$ has a Lipschitz continuous gradient with constant $\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu}$.
Proof.
From Danskins’s theorem [3] one can obtain $f_{\mu}(\cdot)$ is differentiable and $\nabla f_{\mu}(x)=\nabla_{x}\mathcal{L}(x,y^{*}_{\mu}(x))$. Therefore, we have
$\left\|\nabla f_{\mu}(x)-\nabla f_{\mu}(x^{\prime})\right\|=\left\|\nabla_{x}\mathcal{L}(x,y^{*}_{\mu}(x))-\nabla_{x}\mathcal{L}(x^{\prime},y^{*}_{\mu}(x^{\prime}))\right\|$ (30)
Using Lemma 2.2-(II) we can show
$\left\|\nabla f_{\mu}(x)-\nabla f_{\mu}(x^{\prime})\right\|\leq\mathbf{L}_{\mathcal{L}_{1}}\left\|x-x^{\prime}\right\|+\mathbf{L}_{\mathcal{L}_{2}}\left\|y^{*}_{\mu}(x)-y^{*}_{\mu}(x^{\prime})\right\|$ (31)
Considering the result in Lemma 4.3 we have
$\left\|\nabla f_{\mu}(x)-\nabla f_{\mu}(x^{\prime})\right\|\leq\mathbf{L}_{\mathcal{L}_{1}}\left\|x-x^{\prime}\right\|+\mathbf{L}_{\mathcal{L}_{2}}\frac{L_{y\mu}}{\mu}\left\|x-x^{\prime}\right\|$ (32)
$\left\|\nabla f_{\mu}(x)-\nabla f_{\mu}(x^{\prime})\right\|\leq(\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu})\left\|x-x^{\prime}\right\|$ (33)
Next, we aim to evaluate the accuracy of the iterative solution $y_{k}$ in approximating the optimal solution $y^{*}_{\mu}(x_{k})$ by bounding the error between them.
Lemma 4.5.
Suppose Assumptions 2.1, 2.2 hold. Considering that $\{(x_{k},y_{k})\}_{k\geq 0}$ be the sequence generated by the Algorithm. Let $\sigma_{k}=\sigma\leq\frac{2}{L_{yy}^{\Phi}+2\mu}$. Then for any $k\geq 0$,
$\left\|y_{k}-y^{*}_{\mu}(x_{k})\right\|$ $\leq\rho_{D}\|y_{k-1}-y^{*}_{\mu}(x_{k})\|+\sigma L^{\Phi}_{y\theta}\Big{(}\beta^{k}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\gamma\mathbf{L}_{\theta}\sum_{i=0}^{k-1}\beta^{k-i}\left\|s_{i}-x_{i}\right\|\Big{)},$ (34)
where $\rho_{D}\triangleq L_{yy}^{\Phi}/(L_{yy}^{\Phi}+2\mu)$. Moreover, when $L_{yy}^{\Phi}=0$, selecting $\sigma_{k}=\sigma=\frac{1}{\mu}$ and $y_{0}=\mathbf{0}$ we have that for any $k\geq 0$,
$\left\|y_{k}-y^{*}_{\mu}(x_{k})\right\|\leq\frac{L^{\Phi}_{y\theta}}{\mu}\Big{(}\beta^{k}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\gamma\mathbf{L}_{\theta}\sum_{i=0}^{k-1}\beta^{k-i}\left\|s_{i}-x_{i}\right\|\Big{)}.$ (35)
Proof.
From the optimality condition we have that $y^{*}_{\mu}(x)=\mathcal{P}_{\mathcal{Y}}\Big{(}y^{*}_{\mu}(x)+\sigma\nabla_{y}\mathcal{L}_{\mu}(x,y^{*}_{\mu}(x))\Big{)}$ for any $x\in\mathcal{X}$. Moreover, from the update of $y_{k}$ as $y_{k}=\mathcal{P}_{\mathcal{Y}}(y_{k-1}+\sigma G^{y}_{k-1})$ where $G^{y}_{k}\triangleq\nabla_{y}\Phi(x_{k},\theta_{k},y_{k})-\mu(y_{k}-y_{0})$ for any $k\geq 0$, and the non-expansivity of the projection operator, by adding and subtracting $\nabla_{y}\mathcal{L}_{\mu}(x_{k},y_{k-1})$, we have
$\left\|y_{k}-y^{*}_{\mu}(x_{k})\right\|$ $\leq\left\|y_{k-1}+\sigma\Big{(}G^{y}_{k-1}+\nabla_{y}\mathcal{L}_{\mu}(x_{k},y_{k-1})-\nabla_{y}\mathcal{L}_{\mu}(x_{k},y_{k-1})\Big{)}\right.$
$\quad-\left.\Big{(}y^{*}_{\mu}(x_{k})+\sigma\nabla_{y}\mathcal{L}_{\mu}(x,y^{*}_{\mu}(x_{k}))\Big{)}\right\|$
$\leq\|\nabla g_{k}(y_{k-1})-\nabla g_{k}(y^{*}_{\mu}(x_{k}))\|+\sigma\|G_{k-1}^{y}-\nabla_{y}\mathcal{L}_{\mu}(x_{k},y_{k-1})\|$
$\leq\rho_{D}\|y_{k-1}-y^{*}_{\mu}(x_{k})\|+\sigma\|G_{k-1}^{y}-\nabla_{y}\mathcal{L}_{\mu}(x_{k},y_{k-1})\|,$ (36)
where $g_{k}(y)\triangleq\frac{1}{2}\|y\|^{2}+\sigma\mathcal{L}_{\mu}(x_{k},y)$ and $\rho_{D}\triangleq\max\{|1-\sigma(L^{\Phi}_{yy}+\mu)|,|1-\sigma\mu|\}=L_{yy}^{\Phi}/(L_{yy}^{\Phi}+2\mu)$. Next, using the definition of $G^{y}_{k}$ and $\nabla_{y}\mathcal{L}$ and considering Assumption 2.1-(3) we obtain
$\left\|y_{k}-y^{*}_{\mu}(x_{k})\right\|$ $\leq\rho_{D}\|y_{k-1}-y^{*}_{\mu}(x_{k})\|+\sigma\|\nabla_{y}\Phi(x_{k},\theta_{k},y_{k-1})-\nabla_{y}\Phi(x_{k},\theta^{*}(x_{k}),y_{k-1})\|$
$\leq\rho_{D}\|y_{k-1}-y^{*}_{\mu}(x_{k})\|+\sigma L^{\Phi}_{y\theta}\|\theta_{k}-\theta^{*}(x_{k})\|.$ (37)
Using Lemma 4.1 we can show
$\left\|y_{k}-y^{*}_{\mu}(x_{k})\right\|$ $\leq\rho_{D}\|y_{k-1}-y^{*}_{\mu}(x_{k})\|+\sigma L^{\Phi}_{y\theta}(\beta^{k}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\mathbf{L}_{\theta}\sum_{i=0}^{k-1}\beta^{k-i}\left\|x_{i}-x_{i+1}\right\|).$ (38)
Considering $x_{k+1}-x_{k}=\gamma(s_{k}-x_{k})$ we can obtain the desired result.
Lemma 4.6.
Under the premises of Lemma 4.5, for any $k\geq 0$ we have
$\left\|y_{k}-y^{*}_{\mu}(x_{k})\right\|$ $\leq\rho_{D}^{k}\left\|y_{0}-y^{*}_{\mu}(x_{0})\right\|+\frac{\gamma L_{y\mu}}{\mu}\sum_{i=0}^{k-1}\rho_{D}^{k-i}\left\|s_{i}-x_{i}\right\|+\sigma L_{y\theta}^{\Phi}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|\sum_{i=1}^{k}\rho^{k-i}_{D}\beta^{i}$
$\quad+\sigma\gamma L_{y\theta}^{\Phi}\mathbf{L}_{\theta}\sum_{i=1}^{k}\rho^{k-i}_{D}\sum_{j=0}^{i-1}\beta^{i-j}\left\|s_{j}-x_{j}\right\|.$ (39)
Moreover, for any $K\geq 1$,
$\sum_{k=0}^{K-1}\left\|y_{k}-y^{*}_{\mu}(x_{k})\right\|\|s_{k}-x_{k}\|\leq\Gamma_{1}+\Gamma_{2}\sum_{k=0}^{K-1}\|s_{k}-x_{k}\|^{2},$ (40)
where $\Gamma_{1}\triangleq\frac{1}{1-\rho_{D}}\left\|y_{0}-y^{*}_{\mu}(x_{0})\right\|D_{\mathcal{X}}+\frac{\sigma L_{y\theta}^{\Phi}\beta D_{\mathcal{X}}}{(1-\rho_{D})(1-\beta)}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|=\mathcal{O}(\frac{\kappa_{g}}{\mu}L_{y\theta}^{\Phi})$ and $\Gamma_{2}\triangleq\frac{\gamma L_{y\mu}}{\mu(1-\rho_{D})}+\frac{\sigma\gamma\beta L_{y\theta}^{\Phi}\mathbf{L}_{\theta}}{(1-\beta)(1-\rho_{D})}=\mathcal{O}\left(\frac{\kappa_{g}L_{y\theta}^{\Phi}+L_{yx}^{\Phi}}{\mu^{2}}\gamma+\frac{\kappa_{g}^{2}L_{y\theta}^{\Phi}}{\mu}\gamma\right)$.
Proof.
From Lemma 4.5 we have
$\left\|y_{k}-y^{*}_{\mu}(x_{k})\right\|\leq\rho_{D}\left\|y_{k-1}-y^{*}_{\mu}(x_{k})\right\|+\sigma L_{y\theta}^{\Phi}(\beta^{k}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\gamma\mathbf{L}_{\theta}\sum_{i=0}^{k-1}\beta^{k-i}\left\|s_{i}-x_{i}\right\|).$
By adding and subtracting $y^{*}_{\mu}(x_{k-1})$ and using triangle inequality we obtain
$\left\|y_{k}-y^{*}_{\mu}(x_{k})\right\|$ $\leq\rho_{D}(\left\|y_{k-1}-y^{*}_{\mu}(x_{k-1})\right\|+\left\|y^{*}_{\mu}(x_{k})-y^{*}_{\mu}(x_{k-1})\right\|)+\sigma L_{y\theta}^{\Phi}\Big{(}\beta^{k}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|$
$\quad+\gamma\mathbf{L}_{\theta}\sum_{i=0}^{k-1}\beta^{k-i}\left\|s_{i}-x_{i}\right\|\Big{)}$
$\leq\rho_{D}\left\|y_{k-1}-y^{*}_{\mu}(x_{k-1})\right\|+\frac{L_{y\mu}}{\mu}\rho_{D}\left\|x_{k}-x_{k-1}\right\|+\sigma L_{y\theta}^{\Phi}\Big{(}\beta^{k}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|$
$\quad+\gamma\mathbf{L}_{\theta}\sum_{i=0}^{k-1}\beta^{k-i}\left\|s_{i}-x_{i}\right\|\Big{)},$ (41)
where in the last inequality we applied Lemma 4.3.
Continuing the above inequalities recursively we obtain
$\left\|y_{k}-y^{*}_{\mu}(x_{k})\right\|$ $\leq\rho_{D}^{k}\left\|y_{0}-y^{*}_{\mu}(x_{0})\right\|+\frac{\gamma L_{y\mu}}{\mu}\sum_{i=0}^{k-1}\rho_{D}^{k-i}\left\|s_{i}-x_{i}\right\|+\sum_{i=1}^{k}\sigma L_{y\theta}^{\Phi}\Big{(}\rho^{k-i}_{D}\beta^{i}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|$
$\quad+\rho^{k-i}_{D}\gamma\mathbf{L}_{\theta}\sum_{j=0}^{i-1}\beta^{i-j}\left\|s_{j}-x_{j}\right\|\Big{)},$ (42)
which leads to the result in (4.6) by rearranging the terms.
Moreover, using Lemma 8.1 one can verify that for any $K\geq 1$,
$\sum_{k=0}^{K-1}\left\|y_{k}-y^{*}_{\mu}(x_{k})\right\|\|s_{k}-x_{k}\|$ $\leq\frac{1}{1-\rho_{D}}\left\|y_{0}-y^{*}_{\mu}(x_{0})\right\|D_{\mathcal{X}}+\frac{\gamma L_{y\mu}}{\mu(1-\rho_{D})}\sum_{k=0}^{K-1}\|s_{k}-x_{k}\|^{2}$
$\quad+\frac{\sigma L_{y\theta}^{\Phi}\beta D_{\mathcal{X}}}{(1-\rho_{D})(1-\beta)}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\frac{\sigma\gamma\beta L_{y\theta}^{\Phi}\mathbf{L}_{\theta}}{(1-\beta)(1-\rho_{D})}\sum_{k=0}^{K-1}\|s_{k}-x_{k}\|^{2}.$
Finally, note that $\mathbf{L}_{\theta}=\mathcal{O}(\kappa_{g})$, $L_{y\mu}=\mathcal{O}(\kappa_{g})$, $1-\rho_{D}=\mathcal{O}(\mu)$, and $1-\beta=\mathcal{O}(1/\kappa_{g})$.
Lemma 4.7.
Suppose Assumption 2.1, 2.2 hold. For any $k\geq 0$ we have
$\left\|y_{k+1}-y_{k}\right\|$ $\leq 2\rho_{D}^{k}\left\|y_{0}-y^{*}_{\mu}(x_{0})\right\|+\frac{2\gamma L_{y\mu}}{\mu}\sum_{i=0}^{k-1}\rho_{D}^{k-i}\left\|s_{i}-x_{i}\right\|+2\sigma L_{y\theta}^{\Phi}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|\sum_{i=1}^{k}\rho^{k-i}_{D}\beta^{i}$
$\quad+2\sigma\gamma L_{y\theta}^{\Phi}\mathbf{L}_{\theta}\sum_{i=1}^{k}\rho^{k-i}_{D}\sum_{j=0}^{i-1}\beta^{i-j}\left\|s_{j}-x_{j}\right\|+\frac{L_{y\mu}}{\mu}\gamma\left\|s_{k}-x_{k}\right\|.$ (43)
Moreover, for any $K\geq 1$
$\sum_{k=0}^{K-1}\left\|s_{k}-x_{k}\right\|\sum_{i=0}^{k-1}\rho^{k-i}\left\|y_{i+1}-y_{i}\right\|\leq 2\Gamma_{1}+\Gamma_{3}\sum_{k=0}^{K-1}\|s_{k}-x_{k}\|^{2},$ (44)
where $\Gamma_{3}\triangleq\frac{2\Gamma_{2}}{1-\rho}+\frac{L_{y\mu}\gamma}{(1-\rho)\mu}$, and $\Gamma_{1},\Gamma_{2}$ are defined in Lemma 4.6 and $\Gamma_{3}=\mathcal{O}(\frac{\kappa_{g}^{2}L_{y\theta}^{\Phi}+\kappa_{g}L_{yx}^{\Phi}}{\mu^{2}}\gamma+\frac{\kappa_{g}^{3}L_{y\theta}^{\Phi}}{\mu}\gamma)$.
Proof.
Using the triangle inequality one can observe that for any $k\geq 0$,
$\left\|y_{k+1}-y_{k}\right\|\leq\left\|y_{k+1}-y^{*}_{\mu}(x_{k+1})\right\|+\left\|y^{*}_{\mu}(x_{k})-y_{k}\right\|+\left\|y^{*}_{\mu}(x_{k+1})-y^{*}_{\mu}(x_{k})\right\|$ (45)
Using Lemma 4.3 and 4.6 we can obtain
$\left\|y_{k+1}-y_{k}\right\|$ $\leq 2\rho_{D}^{k}\left\|y_{0}-y^{*}_{\mu}(x_{0})\right\|+\frac{2\gamma L_{y\mu}}{\mu}\sum_{i=0}^{k-1}\rho_{D}^{k-i}\left\|s_{i}-x_{i}\right\|+2\sigma L_{y\theta}^{\Phi}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|\sum_{i=1}^{k}\rho^{k-i}_{D}\beta^{i}$
$\quad+2\sigma\gamma L_{y\theta}^{\Phi}\mathbf{L}_{\theta}\sum_{i=1}^{k}\rho^{k-i}_{D}\sum_{j=0}^{i-1}\beta^{i-j}\left\|s_{j}-x_{j}\right\|+\frac{L_{y\mu}}{\mu}\gamma\left\|s_{k}-x_{k}\right\|.$
Moreover, the second part of the lemma can be proved with a similar discussion as in the proof of Lemma 4.6 by employing the results in Lemma 8.1.
4.3 Analysis of Gradients Estimation
In this section, we aim to assess how accurately the gradient estimates, denoted as $G^{x}_{k}$ and $G^{y}_{k}$, approximate the true gradients $\nabla_{x}\mathcal{L}(x_{k},y_{k})$ and $\nabla_{y}\mathcal{L}(x_{k},y_{k})$, respectively. Through Lemmas 4.8 and 4.9, we analyze the bounds on the error in estimating these gradients. We demonstrate that the errors may depend linearly on certain terms or involve consecutive terms, helping us in controlling the error. Later, in the convergence analysis, we will also specify the step-sizes to ensure convergence.
Lemma 4.8.
Suppose Assumptions 2.1 and 2.2 hold. Moreover, let $\{x_{k},\theta_{k},w_{k}\}_{k\geq 0}$ be the sequence generated by Algorithm 1 or 2. Then, for any $k\geq 0$
$\left\|\nabla_{x}\mathcal{L}(x_{k},y_{k})-G^{x}_{k}\right\|$ $\leq\tilde{C}_{2}\beta^{k}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\tilde{C}_{2}\gamma\mathbf{L}_{\theta}\sum_{i=0}^{k-1}\beta^{k-i}\left\|s_{i}-x_{i}\right\|$
$\quad+C_{\theta x}^{g}\Big{(}\rho^{k+1}\|w_{0}-v(x_{0},y_{0})\|+\gamma\mathbf{C}_{v1}\sum_{i=0}^{k-1}\rho^{k-i}\|s_{i}-x_{i}\|$
$\quad+\mathbf{C}_{v2}\sum_{i=0}^{k-1}\rho^{k-i}\|y_{i+1}-y_{i}\|+\tilde{C}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|k\rho^{k}$
$\quad+\gamma\tilde{C}\mathbf{L}_{\theta}\sum_{i=0}^{k}\sum_{j=0}^{i-1}\rho^{k-i}\beta^{i-j}\left\|s_{j}-x_{j}\right\|\Big{)},$ (46)
where $\tilde{C}_{2}\triangleq\frac{C_{\theta}^{\Phi}}{\mu_{g}}L_{\theta x}^{g}+L_{x\theta}^{\Phi}$. Moreover, for any $K\geq 1$
$\sum_{k=0}^{K-1}\left\|s_{k}-x_{k}\right\|\left\|\nabla_{x}\mathcal{L}(x_{k},y_{k})-G^{x}_{k}\right\|\leq\Gamma_{4}+2\mathbf{C}_{v2}\Gamma_{1}+(\mathbf{C}_{v2}\Gamma_{3}+\Gamma_{5})\sum_{k=0}^{K-1}\left\|s_{k}-x_{k}\right\|^{2},$ (47)
where $\Gamma_{1},\Gamma_{3}$ are defined in Lemma 4.6 and 4.7, respectively, $\Gamma_{4}\triangleq D_{\mathcal{X}}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|(\frac{\tilde{C}_{2}}{1-\beta}+\frac{\rho\tilde{C}}{(1-\rho)^{2}})+\frac{\rho C_{\theta x}^{g}D_{\mathcal{X}}}{1-\rho}\left\|w_{0}-v(x_{0},y_{0})\right\|$ and $\Gamma_{5}\triangleq\frac{\gamma\beta\mathbf{L}_{\theta}}{1-\beta}(1+\frac{\tilde{C}}{1-\rho})+\frac{\rho\gamma\mathbf{C}_{v1}C_{\theta x}^{g}}{1-\rho}$. In particular, $\Gamma_{4}+2\mathbf{C}_{v2}\Gamma_{1}=\mathcal{O}(\kappa_{g}^{3}+L_{y\theta}^{\Phi}\kappa_{g}^{2}/\mu)$ and $\mathbf{C}_{v2}\Gamma_{3}+\Gamma_{5}=\mathcal{O}\left(\gamma(\frac{\kappa_{g}^{3}}{\mu^{2}}+\frac{\kappa_{g}^{4}}{\mu})(L_{y\theta}^{\Phi})^{2}+\gamma\frac{\kappa_{g}^{2}}{\mu^{2}}L_{y\theta}^{\Phi}L_{yx}^{\Phi}+\gamma\kappa_{g}^{4}\right)$.
Proof.
We begin the proof by considering the definition of $\nabla_{x}\mathcal{L}(x_{k},y_{k})$ and $G^{x}_{k}$ followed by a triangle inequality to obtain
$\left\|\nabla_{x}\mathcal{L}(x_{k},y_{k})-G^{x}_{k}\right\|$ $=\left\|\nabla_{x}\Phi(x_{k},\theta^{*}(x_{k}),y_{k})-\nabla_{\theta x}^{2}g(x_{k},\theta^{*}(x_{k}))v(x_{k},y_{k})\right.$
$\quad-\left.[\nabla_{x}\Phi(x_{k},\theta_{k},y_{k})-\nabla_{\theta x}^{2}g(x_{k},\theta_{k})w_{k+1}]\right\|$
$\leq\left\|\nabla_{x}\Phi(x_{k},\theta^{*}(x_{k}),y_{k})-\nabla_{x}\Phi(x_{k},\theta_{k},y_{k})\right\|+\left\|\nabla_{\theta x}^{2}g(x_{k},\theta_{k})w_{k+1}\right.$
$\quad-\left.\nabla_{\theta x}^{2}g(x_{k},\theta^{*}(x_{k}))v(x_{k},y_{k})\right\|$ (48)
Combining Assumption 2.1-(1) together with adding and subtracting $\nabla_{\theta x}^{2}g(x_{k},\theta_{k})v(x_{k},y_{k})$ to the second term of RHS lead to
$\left\|\nabla_{x}\mathcal{L}(x_{k},y_{k})-G^{x}_{k}\right\|$ $\leq L_{x\theta}^{\Phi}\left\|\theta_{k}-\theta^{*}(x_{k})\right\|+\left\|\nabla_{\theta x}^{2}g(x_{k},\theta_{k})[w_{k+1}-v(x_{k},y_{k})]\right.$
$\quad+\left.[\nabla_{\theta x}^{2}g(x_{k},\theta_{k})-\nabla_{\theta x}^{2}g(x_{k},\theta^{*}(x_{k}))]v(x_{k},y_{k})\right\|$
$\leq L_{x\theta}^{\Phi}\left\|\theta_{k}-\theta^{*}(x_{k})\right\|+C_{\theta x}^{g}\left\|w_{k+1}-v(x_{k},y_{k})\right\|+\frac{C_{\theta}^{\Phi}}{\mu_{g}}L_{\theta x}^{g}\left\|\ \theta_{k}-\theta^{*}(x_{k})\right\|$ (49)
where the last inequality is obtained using Assumption 2.2 and the triangle inequality. Next, utilizing Lemma 4.1 and 4.2 we can further provide upper-bounds for the terms in RHS of the above inequality to obtain the desired result in (4.8).
To show the result in (47), note that $\sum_{k=0}^{K-1}k\rho^{k}\leq\rho/(1-\rho)^{2}$. Moreover, one can multiply $\left\|s_{k}-x_{k}\right\|$ to each term on the RHS of (4.8) and sum over $k$ from 00 to $K-1$ and use Lemma 8.1 and (44). Finally, note that $\mathbf{C}_{v1}=\mathcal{O}(\kappa_{g}^{3})$, $\mathbf{C}_{v2}=\mathcal{O}(\kappa_{g})$, $\tilde{C}=\mathcal{O}(\kappa_{g})$, $\tilde{C}_{2}=\mathcal{O}(\kappa_{g})$, and $1-\rho=\mathcal{O}(1/\kappa_{g})$.
Lemma 4.9.
Using the Assumption 2.1-(3), we can find a bound of $\left\|\nabla_{y}\mathcal{L}(x_{k},y_{k})-G_{k}^{y}\right\|$ as follows
$\left\|\nabla_{y}\mathcal{L}(x_{k},y_{k})-G_{k}^{y}\right\|\leq\beta^{k}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\gamma L_{y\theta}^{\Phi}\mathbf{L}_{\theta}\sum_{i=0}^{k-1}\beta^{k-i}\left\|s_{i}-x_{i}\right\|+\mu D_{\mathcal{Y}}.$ (50)
Proof.
Following the definition of $\nabla_{y}\mathcal{L}(x_{k},y_{k})$ in (6b) and $G_{k}^{y}$ in line 6 of the algorithms, we have that
$\left\|\nabla_{y}\mathcal{L}(x_{k},y_{k})-G_{k}^{y}\right\|$ $=\left\|\nabla_{y}\Phi(x_{k},\theta^{*}(x_{k}),y_{k})-\nabla_{y}\Phi(x_{k},\theta_{k},y_{k})+\mu(y_{k}-y_{0})\right\|$
$\leq L_{y\theta}^{\Phi}\left\|\theta_{k}-\theta^{*}(x_{k})\right\|+\mu D_{\mathcal{Y}},$ (51)
where in the last inequality we use Lipschitz continuity of $\nabla_{y}\Phi(x,\cdot,y)$ and boundedness of set $Y$. Next, utilizing Lemma 4.1 we can further provide upper-bounds for the term in RHS of the above inequality as follows
$\left\|\nabla_{y}\mathcal{L}(x_{k},y_{k})-G_{k}^{y}\right\|\leq L_{y\theta}^{\Phi}(\beta^{k}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\mathbf{L}_{\theta}\sum_{i=0}^{k-1}\beta^{k-i}\left\|x_{i}-x_{i+1}\right\|)+\mu D_{\mathcal{Y}}.$
The result follows by noting that $\left\|x_{k+1}-x_{k}\right\|=\gamma\left\|s_{k}-x_{k}\right\|$ for any $k\geq 0$.
5 Convergence Results
In this section, we commence by analyzing the complexity of the i-BRPD:OPF method in Subsection 5.1. Later, in Subsection 5.2, we delve into the iteration complexity analysis of the i-BRPD:FP method across various setups.
5.1 Convergence Rate of Algorithm 1
Theorem 5.1.
Suppose Assumptions 2.1, 2.2 hold and let $\{(x_{k},y_{k},\theta_{k})\}_{k\geq 0}$ be the sequence generated by Algorithm 1 with step-sizes $\gamma_{k}=\gamma>0$, $\eta_{k}=\eta>0$, $\alpha>0$, and $\sigma_{k}=\sigma\leq\frac{2}{L^{\Phi}_{yy}+2\mu}$ and regularization parameter $\mu>0$. Then for any $K\geq 1$ we have that
$\frac{1}{K}\sum_{k=0}^{K-1}\mathcal{G}_{\mathcal{X}}(x_{k},y_{k})$ $\leq\frac{f(x_{0})-f(x^{*})}{\gamma K}+\frac{B_{1}}{K}+\gamma B_{2}D_{\mathcal{X}}^{2}+\frac{\mu}{2\gamma K}D_{\mathcal{Y}}^{2},$ (52)
$\frac{1}{K}\sum_{k=0}^{K-1}\mathcal{G}_{\mathcal{Y}}(x_{k},y_{k})$ $\leq\frac{2\left\|y_{0}-y^{*}_{\mu}(x_{0})\right\|}{\sigma(1-\rho_{D})K}+\gamma B_{3}D_{\mathcal{X}}+\frac{3}{(1-\rho_{D})(1-\beta)}\frac{L_{y\theta}^{\Phi}}{K}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\mu D_{\mathcal{Y}},$ (53)
for some constants $B_{1},B_{2},B_{3}>0$ where $B_{1}=\mathcal{O}(\kappa_{g}^{3}+L_{y\theta}^{\Phi}\kappa_{g}^{2}/\mu)$, $B_{2}=\mathcal{O}\big{(}(\frac{\kappa_{g}^{3}}{\mu^{2}}+\frac{\kappa_{g}^{4}}{\mu})(L_{y\theta}^{\Phi})^{2}+\frac{\kappa_{g}^{2}}{\mu^{2}}L_{y\theta}^{\Phi}L_{yx}^{\Phi}+\kappa_{g}^{4}\big{)}$, and $B_{3}\triangleq\mathcal{O}(\frac{\kappa_{g}}{\mu^{2}}+\frac{\kappa_{g}^{2}}{\mu})$.
Proof.
From Lipschitz continuity of $\nabla f_{\mu}$ we have that
$f_{\mu}(x_{k+1})$ $\leq f_{\mu}(x_{k})+\langle\nabla f_{\mu}(x_{k}),x_{k+1}-x_{k}\rangle+\frac{1}{2}(\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu})\|{x_{k+1}-x_{k}}\|^{2}$
$=f_{\mu}(x_{k})+\langle\nabla f_{\mu}(x_{k})-\nabla_{x}\mathcal{L}(x_{k},y_{k}),x_{k+1}-x_{k}\rangle+\langle\nabla_{x}\mathcal{L}(x_{k},y_{k})-G_{k}^{x},x_{k+1}-x_{k}\rangle$
$\quad+\langle G_{k}^{x},x_{k+1}-x_{k}\rangle+\frac{1}{2}(\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu})\|{x_{k+1}-x_{k}}\|^{2},$ (54)
where the last equality follows from adding and subtracting the terms $\langle G_{k}^{x}+\nabla_{x}\mathcal{L}(x_{k},y_{k}),x_{k+1}-x_{k}\rangle$ to RHS. Recall that $x_{k+1}=\gamma s_{k}+(1-\gamma)x_{k}$. Moreover, we define $s^{\prime}_{k}\triangleq\operatorname*{\arg\!\max}_{s\in X}\langle\nabla_{x}\mathcal{L}(x_{k},y_{k}),x_{k}-s\rangle$ then we have that
$\langle G_{k}^{x},x_{k+1}-x_{k}\rangle$ $=\gamma\langle G_{k}^{x},s_{k}-x_{k}\rangle$
$\leq\gamma\langle G_{k}^{x},s^{\prime}_{k}-x_{k}\rangle$
$=\gamma\langle\nabla_{x}\mathcal{L}(x_{k},y_{k})_{k},s^{\prime}_{k}-x_{k}\rangle+\gamma\langle G_{k}^{x}-\nabla_{x}\mathcal{L}(x_{k},y_{k}),s^{\prime}_{k}-x_{k}\rangle.$ (55)
Next, using (5.1) within (5.1), rearranging the terms, and using Cauchy Schwartz inequality imply that
$\gamma\langle\nabla_{x}\mathcal{L}(x_{k},y_{k}),x_{k}-s^{\prime}_{k}\rangle$ $\leq f_{\mu}(x_{k})-f_{\mu}(x_{k+1})+\|\nabla f_{\mu}(x_{k})-\nabla_{x}\mathcal{L}(x_{k},y_{k})\|\|x_{k+1}-x_{k}\|$
$\quad+\gamma\|\nabla_{x}\mathcal{L}(x_{k},y_{k})-G_{k}^{x}\|\|s_{k}-s^{\prime}_{k}\|$
$\quad+\frac{1}{2}(\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu})\|{x_{k+1}-x_{k}}\|^{2}.$ (56)
From Lemma 4.4 we know that $\nabla f_{\mu}(x)=\nabla_{x}\mathcal{L}(x,y^{*}_{\mu}(x))$. Therefore, using Lipschitz continuity of $\mathcal{L}(x,\cdot)$, $\left\|s^{\prime}_{k}-x_{k}\right\|\leq D_{\mathcal{X}}$, and $\|{x_{k+1}-x_{k}}\|=\gamma\left\|s_{k}-x_{k}\right\|$ we obtain
$\gamma\langle\nabla_{x}\mathcal{L}(x_{k},y_{k}),x_{k}-s^{\prime}_{k}\rangle$ $\leq f_{\mu}(x_{k})-f_{\mu}(x_{k+1})+\mathbf{L}_{\mathcal{L}_{2}}\gamma\left\|y_{k}-y_{\mu}^{*}(x_{k})\right\|\left\|s_{k}-x_{k}\right\|$
$\quad+\gamma D_{\mathcal{X}}\|\nabla_{x}\mathcal{L}(x_{k},y_{k})-G_{k}^{x}\|+\frac{1}{2}(\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu})\gamma^{2}\left\|s_{k}-x_{k}\right\|^{2}.$ (57)
Summing the above inequality over $k=0$ to $K-1$ dividing both sides by $\gamma K$, and recalling $\mathcal{G}_{\mathcal{X}}(x_{k},y_{k})=\sup_{s\in\mathcal{X}}\langle\nabla_{x}\mathcal{L}(x_{k},y_{k}),x_{k}-s\rangle$ imply that for any $K\geq 1$,
$\frac{1}{K}\sum_{k\in\mathcal{K}}\mathcal{G}_{\mathcal{X}}(x_{k},y_{k})$ $\leq\frac{(f_{\mu}(x_{0})-f_{\mu}(x_{K}))}{\gamma K}+\frac{\mathbf{L}_{\mathcal{L}_{2}}}{K}\sum_{k=0}^{K-1}\left\|s_{k}-x_{k}\right\|\left\|y_{k}-y^{*}_{\mu}(x_{k})\right\|$
$\quad+\frac{D_{\mathcal{X}}}{K}\sum_{k=0}^{K-1}\|\nabla_{x}\mathcal{L}(x_{k},y_{k})-G_{k}^{x}\|+\gamma(\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu})\frac{1}{K}\sum_{k=0}^{K-1}\left\|s_{k}-x_{k}\right\|^{2}.$
By noting that $\left\|s_{k}-x_{k}\right\|\leq D_{\mathcal{X}}$ due to the boundedness of set $\mathcal{X}$, we can apply Lemmas 4.6, 4.8 to the above inequality to obtain
$\frac{1}{K}\sum_{k=0}^{K-1}\mathcal{G}_{\mathcal{X}}(x_{k},y_{k})$ $\leq\frac{(f_{\mu}(x_{0})-f_{\mu}(x_{K}))}{\gamma K}+\frac{\mathbf{L}_{\mathcal{L}_{2}}}{K}\left(\Gamma_{1}+\Gamma_{2}D_{\mathcal{X}}^{2}K\right)$
$\quad+\frac{\Gamma_{4}+2\mathbf{C}_{v2}\Gamma_{1}}{K}D_{\mathcal{X}}+(\mathbf{C}_{v2}\Gamma_{3}+\Gamma_{5})D_{\mathcal{X}}^{2}+\gamma\left(\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu}\right)D_{\mathcal{X}}^{2}.$
From the definition of $f_{\mu}$ and boundedness of $\mathcal{Y}$ we have $f_{\mu}(x_{0})-f_{\mu}(x_{K})\leq f(x_{0})-f(x^{*})+\frac{\mu}{2}D_{\mathcal{Y}}^{2}$. Therefore, for any $K>2$
$\frac{1}{K}\sum_{k=0}^{K-1}\mathcal{G}_{\mathcal{X}}(x_{k},y_{k})$ $\leq\frac{(f(x_{0})-f(x^{*}))}{\gamma K}+\frac{\mathbf{L}_{\mathcal{L}_{2}}\Gamma_{1}+D_{\mathcal{X}}(\Gamma_{4}+2\mathbf{C}_{v2}\Gamma_{1})}{K}$
$\quad+(\mathbf{L}_{\mathcal{L}_{2}}\Gamma_{2}+\mathbf{C}_{v2}\Gamma_{3}+\Gamma_{5})D_{\mathcal{X}}^{2}+\gamma\left(\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu}\right)D_{\mathcal{X}}^{2}+\frac{\mu}{2\gamma K}D_{\mathcal{Y}}^{2}.$
Finally, recall that $\Gamma_{1}=\mathcal{O}(\frac{\kappa_{g}}{\mu}L_{y\theta}^{\Phi})$, $\Gamma_{2}=\mathcal{O}(\frac{\kappa_{g}L_{y\theta}^{\Phi}+L_{yx}^{\Phi}}{\mu^{2}}\gamma+\frac{\kappa_{g}^{2}L_{y\theta}^{\Phi}}{\mu}\gamma)$, $\Gamma_{3}=\mathcal{O}(\frac{\kappa_{g}^{2}L_{y\theta}^{\Phi}+\kappa_{g}L_{yx}^{\Phi}}{\mu^{2}}\gamma+\frac{\kappa_{g}^{3}L_{y\theta}^{\Phi}}{\mu}\gamma)$, $\Gamma_{4}+2\mathbf{C}_{v2}\Gamma_{1}=\mathcal{O}(\kappa_{g}^{3}+L_{y\theta}^{\Phi}\kappa_{g}^{2}/\mu)$, $\mathbf{C}_{v2}\Gamma_{3}+\Gamma_{5}=\mathcal{O}((\frac{\kappa_{g}^{3}}{\mu^{2}}+\frac{\kappa_{g}^{4}}{\mu})(L_{y\theta}^{\Phi})^{2}\gamma+\frac{\kappa_{g}^{2}}{\mu^{2}}L_{y\theta}^{\Phi}L_{yx}^{\Phi}\gamma+\kappa_{g}^{4}\gamma)$, $\mathbf{L}_{\mathcal{L}_{1}}=\mathcal{O}(\kappa_{g}^{3})$, and $\mathbf{L}_{\mathcal{L}_{2}}=\mathcal{O}(\kappa_{g})$. Putting the pieces together the desired result can be obtained.
Next, we obtain an upper-bound for the dual gap function $\mathcal{G}_{\mathcal{Y}}(x_{k},y_{k})=\frac{1}{\sigma}\left\|y_{k}-\mathcal{P}_{\mathcal{Y}}(y_{k}+\sigma\nabla_{y}\mathcal{L}(x_{k},y_{k}))\right\|$. In particular, using the triangle inequality and by adding and subtracting $\mathcal{P}_{\mathcal{Y}}\big{(}y_{k}+\sigma G_{k}^{y}\big{)}$, we have that for any $k\geq 0$,
$\mathcal{G}_{\mathcal{Y}}(x_{k},y_{k})\leq\frac{1}{\sigma}\big{(}\left\|y_{k}-\mathcal{P}_{\mathcal{Y}}\big{(}y_{k}+\sigma G_{k}^{y}\big{)}\right\|+\left\|\mathcal{P}_{\mathcal{Y}}\big{(}y_{k}+\sigma G_{k}^{y}\big{)}-\mathcal{P}_{\mathcal{Y}}(y_{k}+\sigma\nabla_{y}\mathcal{L}(x_{k},y_{k}))\right\|\big{)}.$ (58)
Considering non-expansivity of the projection mapping we have
$\mathcal{G}_{\mathcal{Y}}(x_{k},y_{k})\leq\frac{1}{\sigma}(\left\|y_{k}-y_{k+1}\right\|+\sigma\left\|G_{k}^{y}-\nabla_{y}\mathcal{L}(x_{k},y_{k})\right\|).$ (59)
Based on the conclusion of Lemma 4.9 we can get
$\mathcal{G}_{\mathcal{Y}}(x_{k},y_{k})\leq\frac{1}{\sigma}\left\|y_{k}-y_{k+1}\right\|+L_{y\theta}^{\Phi}\beta^{k}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\gamma L_{y\theta}^{\Phi}\mathbf{L}_{\theta}\sum_{i=0}^{k-1}\beta^{k-i}\left\|s_{i}-x_{i}\right\|+\mu D_{\mathcal{Y}}.$ (60)
Moreover, using Lemma 4.7 and rearragning the terms we conclude that for any $k\geq 0$,
$\mathcal{G}_{\mathcal{Y}}(x_{k},y_{k})$ $\leq\frac{2}{\sigma}\rho_{D}^{k}\left\|y_{0}-y^{*}_{\mu}(x_{0})\right\|+\frac{2\gamma L_{y\mu}}{\sigma\mu}\sum_{i=0}^{k-1}\rho_{D}^{k-i}\left\|s_{i}-x_{i}\right\|+2L_{y\theta}^{\Phi}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|\sum_{i=1}^{k}\rho^{k-i}_{D}\beta^{i}$
$\quad+2\gamma L_{y\theta}^{\Phi}\mathbf{L}_{\theta}\sum_{i=1}^{k}\rho^{k-i}_{D}\sum_{j=0}^{i-1}\beta^{i-j}\left\|s_{j}-x_{j}\right\|+\frac{L_{y\mu}}{\mu\sigma}\gamma\left\|s_{k}-x_{k}\right\|$
$\quad+L_{y\theta}^{\Phi}\beta^{k}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\gamma L_{y\theta}^{\Phi}\mathbf{L}_{\theta}\sum_{i=0}^{k-1}\beta^{k-i}\left\|s_{k}-x_{k}\right\|+\mu D_{\mathcal{Y}}.$ (61)
Summing the above inequality over $k=0$ to $K-1$, using Lemma 8.1 together with boundedness of set $\mathcal{X}$ lead to
$\frac{1}{K}\sum_{k=0}^{K-1}\mathcal{G}_{\mathcal{Y}}(x_{k},y_{k})$ $\leq\frac{2\left\|y_{0}-y^{*}_{\mu}(x_{0})\right\|}{\sigma(1-\rho_{D})K}+\frac{2L_{y\mu}\rho_{D}\gamma D_{\mathcal{X}}}{\mu(1-\rho_{D})\sigma}+\frac{2L_{y\theta}^{\Phi}\mathbf{L}_{\theta}\beta\gamma D_{\mathcal{X}}}{(1-\beta)(1-\rho_{D})}+\frac{L_{y\mu}}{\mu\sigma}\gamma D_{\mathcal{X}}$
$\quad+\frac{L_{y\theta}^{\Phi}\mathbf{L}_{\theta}\beta\gamma D_{\mathcal{X}}}{1-\beta}+\frac{2L_{y\theta}^{\Phi}}{(1-\rho_{D})(1-\beta)K}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\mu D_{\mathcal{Y}}.$ (62)
Therefore, the desired result can be concluded by noting that $1-\rho_{D}=\mathcal{O}(\mu)$, $1-\beta=\mathcal{O}(1/\kappa_{g})$, and $\mathbf{L}_{\theta}=\mathcal{O}(\kappa_{g})$.
Corollary 5.2.
Under the premises of Theorem 5.1, let $\mu=\mathcal{O}(\epsilon)$ and $\gamma=\mathcal{O}(\epsilon^{3})$, then for any $K\geq 1$, there exists $t\in\{0,\ldots,K-1\}$ such that $(x_{t},y_{t})\in\mathcal{X}\times\mathcal{Y}$ satisfies $\mathcal{G}_{\mathcal{Z}}(x_{t},y_{t})\leq\epsilon$ within $\mathcal{O}(\frac{\kappa_{g}^{3}}{\epsilon^{4}}+\frac{\kappa_{g}^{5}}{\epsilon^{2}})$ iterations.
Proof.
Summing up the results on the primal and dual gap functions in (52) and (53) and defining $k^{*}\triangleq\operatorname*{\arg\!\min}_{k=0}^{K-1}\mathcal{G}_{\mathcal{Z}}(x_{k},y_{k})$ implies that $\mathcal{G}_{\mathcal{Z}}(x_{k^{*}},y_{k^{*}})\leq\mathcal{O}(\frac{1}{\gamma K}+\gamma B_{2}+\mu)$ where $B_{2}=\mathcal{O}(\frac{\kappa_{g}^{3}}{\mu^{2}}+\frac{\kappa_{g}^{4}}{\mu})$. Selecting $\mu=\kappa_{g}^{3/4}/K^{1/4}$ and $\gamma=1/(\kappa_{g}K)^{3/4}$ implies that $\mathcal{G}_{\mathcal{Z}}(x_{k^{*}},y_{k^{*}})\leq\mathcal{O}(\frac{\kappa_{g}^{3/4}}{K^{1/4}}+\frac{\kappa_{g}^{2.5}}{\sqrt{K}})$. Therefore, to achieve $\mathcal{G}_{\mathcal{Z}}(x_{k^{*}},y_{k^{*}})\leq\epsilon$, Algorithm 1 requires $K=\mathcal{O}(\frac{\kappa_{g}^{3}}{\epsilon^{4}}+\frac{\kappa_{g}^{5}}{\epsilon^{2}})$ iterations.
Corollary 5.3.
Suppose the premises of Theorem 5.1 holds and we further assume that $L_{yy}^{\Phi}=0$. Let $\mu=\mathcal{O}(\epsilon)$, $\gamma=\mathcal{O}(\epsilon^{3})$ then for any $K\geq 1$, there exists $t\in\{0,\ldots,K-1\}$ such that $(x_{t},y_{t})\in\mathcal{X}\times\mathcal{Y}$ satisfies $\mathcal{G}_{\mathcal{Z}}(x_{t},y_{t})\leq\epsilon$ within $\mathcal{O}(\frac{\kappa_{g}^{4}}{\epsilon^{3}})$ iterations.
Proof.
Assuming $L^{\Phi}_{yy}=0$, then, $\rho_{D}=0$ and $\sigma=\mathcal{O}(\frac{1}{\mu})$. By revisiting Lemmas 4.5, 4.6, and 4.7 we obtain $\Gamma_{1}=\mathcal{O}(\frac{\kappa_{g}}{\mu}L_{y\theta}^{\Phi})$, $\Gamma_{2}=\mathcal{O}\left(\frac{\kappa_{g}L_{y\theta}^{\Phi}+L_{yx}^{\Phi}}{\mu}\gamma+\frac{\kappa_{g}^{2}L_{y\theta}^{\Phi}}{\mu}\gamma\right)$, and $\Gamma_{3}=\mathcal{O}(\frac{\kappa_{g}^{2}L_{y\theta}^{\Phi}+\kappa_{g}L_{yx}^{\Phi}}{\mu}\gamma+\frac{\kappa_{g}^{3}L_{y\theta}^{\Phi}}{\mu}\gamma)$. Similar to the proof of Corollary 5.2, by updating $B_{2}=\mathcal{O}(\frac{\kappa_{g}^{4}}{\mu})$ and selecting $\mu=\mathcal{O}(1/(K)^{1/3})$, $\gamma=\mathcal{O}(1/(K)^{2/3})$, Algorithm 1 requires $K=\mathcal{O}(\frac{\kappa_{g}^{4}}{\epsilon^{3}})$ iterations to achieve $\mathcal{G}_{\mathcal{Z}}(x_{k^{*}},y_{k^{*}})\leq\epsilon$.
Remark 5.1.
Our results in Theorem 5.1 represent a novel bound for the constrained nonconvex-concave saddle point problem, subject to a strongly convex problem in the lower-level setting. To the best of our knowledge, there has been no exploration of any projection-free algorithm alternative in this setup.
5.2 Convergence Rate of Algorithm 2
Theorem 5.4.
Suppose Assumptions 2.1, 2.2 hold and let $\{(x_{k},y_{k},\theta_{k})\}_{k\geq 0}$ be the sequence generated by Algorithm 2 with step-sizes $\gamma_{k}=\gamma>0$, $\eta_{k}=\eta>0$, $\alpha>0$, $\tau_{k}=\tau>0$, and $\sigma_{k}=\sigma\leq\frac{2}{L^{\Phi}_{yy}+2\mu}$ and regularization parameter $\mu>0$. Then, for any $K\geq\bar{K}=\mathcal{O}(\mu\log(\kappa_{g}/\mu))$ we have that
$\frac{1}{\tau^{2}K}\sum_{k=0}^{K-1}\left\|s_{k}-x_{k}\right\|^{2}$ $\leq\frac{2(f(x_{0})-f(x^{*}))+\mu D_{\mathcal{Y}}}{\tau\gamma K}+\frac{\gamma B_{4}}{\tau K},$ (63)
$\frac{1}{K}\sum_{k=0}^{K-1}\mathcal{G}_{\mathcal{Y}}(x_{k},y_{k})$ $\leq\frac{2\left\|y_{0}-y^{*}_{\mu}(x_{0})\right\|}{\sigma(1-\rho_{D})K}+\gamma B_{3}D_{\mathcal{X}}+\frac{3}{(1-\rho_{D})(1-\beta)}\frac{L_{y\theta}^{\Phi}}{K}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\mu D_{\mathcal{Y}},$ (64)
for some constant $B_{3},B_{4}>0$ where $B_{3}$ is defined in Theorem 5.1 and $B_{4}=\mathcal{O}\big{(}(\frac{\kappa_{g}^{3}}{\mu^{2}}+\frac{\kappa_{g}^{4}}{\mu})(L_{y\theta}^{\Phi})^{2}+\frac{\kappa_{g}^{2}}{\mu^{2}}L_{y\theta}^{\Phi}L_{yx}^{\Phi}+\kappa_{g}^{4}\big{)}$.
Proof.
Similar to (5.1), from Lipschitz continuity of $\nabla f_{\mu}$ we have that
$f_{\mu}(x_{k+1})$ $\leq f_{\mu}(x_{k})+\langle\nabla f_{\mu}(x_{k})-\nabla_{x}\mathcal{L}(x_{k},y_{k}),x_{k+1}-x_{k}\rangle+\langle\nabla_{x}\mathcal{L}(x_{k},y_{k})-G_{k}^{x},x_{k+1}-x_{k}\rangle$
$\quad+\langle G_{k}^{x},x_{k+1}-x_{k}\rangle+\frac{1}{2}(\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu})\|{x_{k+1}-x_{k}}\|^{2}.$ (65)
Recall the update of $s_{k}=\mathcal{P}_{\mathcal{X}}(x_{k}-\tau G_{k}^{x})$. The optimality condition of this step implies that
$\left\langle G_{k}^{x},x_{k+1}-x_{k}\right\rangle=\gamma\left\langle G_{k}^{x},s_{k}-x_{k}\right\rangle\leq-\frac{\gamma}{\tau}\left\|s_{k}-x_{k}\right\|^{2}.$ (66)
Combining (66) with (5.2) and using $\|x_{k+1}-x_{k}\|=\gamma\left\|s_{k}-x_{k}\right\|$ we obtain
$\frac{\gamma}{2\tau}\left\|s_{k}-x_{k}\right\|^{2}$ $\leq f_{\mu}(x_{k})-f_{\mu}(x_{k+1})+\gamma\mathbf{L}_{\mathcal{L}_{2}}\left\|y_{k}-y_{\mu}^{*}(x_{k})\right\|\left\|s_{k}-x_{k}\right\|+\gamma\|\nabla_{x}\mathcal{L}(x_{k},y_{k})-G_{k}^{x}\|\left\|s_{k}-x_{k}\right\|$
$\quad+\frac{1}{2}(\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu})\gamma^{2}\left\|s_{k}-x_{k}\right\|^{2}-\frac{\gamma}{2\tau}\left\|s_{k}-x_{k}\right\|^{2}.$ (67)
Summing the above inequality over $k=0$ to $K-1$, dividing both sides by $\gamma K$, and using Lemmas 4.6, 4.8 we obtain
$\frac{1}{2\tau K}\sum_{k=0}^{K-1}\left\|s_{k}-x_{k}\right\|^{2}$ $\leq\frac{f_{\mu}(x_{0})-f_{\mu}(x_{K})}{\gamma K}+\frac{\mathbf{L}_{\mathcal{L}_{2}}\Gamma_{1}+2\mathbf{C}_{v2}\Gamma_{1}+\Gamma_{4}}{K}$
$\quad+\left(\mathbf{L}_{\mathcal{L}_{2}}\Gamma_{2}+\mathbf{C}_{v2}\Gamma_{3}+\Gamma_{5}+\frac{1}{2}(\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu})\gamma-\frac{1}{2\tau}\right)\frac{1}{K}\sum_{k=0}^{K-1}\left\|s_{k}-x_{k}\right\|^{2}.$ (68)
Selecting step-size $\tau$ such that $\tau\leq 2/(\mathbf{L}_{\mathcal{L}_{2}}\Gamma_{2}+\mathbf{C}_{v2}\Gamma_{3}+\Gamma_{5}+\frac{1}{2}(\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu})\gamma)$ we can drop the last term in RHS of the above inequality. Therefore, using the fact that $f_{\mu}(x_{0})-f_{\mu}(x_{K})\leq f(x_{0})-f(x^{*})+\frac{\mu}{2}D_{\mathcal{Y}}^{2}$ and multiplying both sides by $2/\tau$ the desired result can be obtained.
Next, using the same line of proof as in Theorem 5.1 we can derive a bound for the dual gap function as in (5.1). Therefore, using boundedness of constraint set $\mathcal{X}$ and by noting that for any $k\geq 2$ $\sum_{i=1}^{k}\rho_{D}^{k-i}\beta^{i}\leq(\frac{\rho_{D}^{k/2}}{1-\beta}+\frac{\beta^{k/2}}{1-\rho_{D}})\beta$, we conclude that for any $k\geq 2$,
$\mathcal{G}_{\mathcal{Y}}(x_{k},y_{k})$ $\leq\frac{2}{\sigma}\rho_{D}^{k}\left\|y_{0}-y^{*}_{\mu}(x_{0})\right\|+\frac{2\gamma L_{y\mu}}{\sigma\mu(1-\rho_{D})}D_{\mathcal{X}}+2L_{y\theta}^{\Phi}\beta\left(\frac{\rho_{D}^{k/2}}{1-\beta}+\frac{\beta^{k/2}}{1-\rho_{D}}\right)\left\|\theta_{0}-\theta^{*}(x_{0})\right\|$
$\quad+\frac{2\gamma\beta L_{y\theta}^{\Phi}\mathbf{L}_{\theta}}{(1-\rho_{D})(1-\beta)}D_{\mathcal{X}}+\frac{L_{y\mu}}{\mu\sigma}\gamma D_{\mathcal{X}}+L_{y\theta}^{\Phi}\beta^{k}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\frac{\gamma L_{y\theta}^{\Phi}\mathbf{L}_{\theta}}{1-\beta}D_{\mathcal{X}}+\mu D_{\mathcal{Y}}.$ (69)
Corollary 5.5.
Under the premises of Theorem 5.4, let $\mu=\mathcal{O}(\epsilon)$, $\gamma=\mathcal{O}(\epsilon^{3})$ and $\tau\leq\mathcal{O}(\epsilon(\kappa_{g}^{2}L_{y\theta}^{\Phi}(\kappa_{g}L_{y\theta}^{\Phi}+L_{yx}^{\Phi}))^{-1})$, then for any $K\geq\bar{K}=\mathcal{O}(\mu\log(\kappa_{g}/\mu))$, there exists $t\in\{\bar{K},\ldots,K-1\}$ such that $(x_{t},y_{t})\in\mathcal{X}\times\mathcal{Y}$ satisfies $\mathcal{G}_{\mathcal{Z}}(x_{t},y_{t})\leq\epsilon$ within $\mathcal{O}(\frac{\kappa_{g}^{3}}{\epsilon^{5}}\log(\frac{\kappa_{g}}{\epsilon})+\frac{\kappa_{g}^{8}}{\epsilon^{2.5}}+\frac{\kappa_{g}^{6}}{\epsilon^{2}})$ iterations.
Proof.
Considering the result in (5.2) of Theorem 5.4 and multiplying both sides by $K/(K-\bar{K})$ for some $0<\bar{K}<K$. Then, we have that for any $K\geq 2$,
$\frac{1}{\tau^{2}(K-\bar{K})}\sum_{k=\bar{K}}^{K-1}\left\|s_{k}-x_{k}\right\|^{2}$ $\leq\frac{1}{\tau^{2}(K-\bar{K})}\sum_{k=0}^{K-1}\left\|s_{k}-x_{k}\right\|^{2}$
$\leq\frac{2(f(x_{0})-f(x^{*}))+\mu D_{\mathcal{Y}}^{2}}{\tau\gamma(K-\bar{K})}+\frac{2(\mathbf{L}_{\mathcal{L}_{2}}\Gamma_{1}+2\mathbf{C}_{v2}\Gamma_{1}+\Gamma_{4})}{\tau(K-\bar{K})}$
Let $k^{*}=\operatorname*{\arg\!\min}_{k=\bar{K}}^{K-1}\left\|s_{k}-x_{k}\right\|$. Lower-bounding the LHS and taking the square root from both sides of the above inequality, we obtain
$\mathcal{G}_{\mathcal{X}}(x_{k^{*}},y_{k^{*}})=\frac{1}{\tau}\left\|s_{k^{*}}-x_{k^{*}}\right\|$ $\leq\sqrt{\frac{2(f(x_{0})-f(x^{*}))+\mu D_{\mathcal{Y}}^{2}}{\gamma\tau(K-\bar{K})}}+\sqrt{\frac{2(\mathbf{L}_{\mathcal{L}_{2}}\Gamma_{1}+2\mathbf{C}_{v2}\Gamma_{1}+\Gamma_{4})}{\tau(K-\bar{K})}}$
$=\mathcal{O}\left(\frac{1}{\sqrt{\gamma\tau K}}+\sqrt{\frac{\kappa_{g}^{3}+L_{y\theta}^{\Phi}\kappa_{g}^{2}/\mu}{\tau K}}\right),$ (70)
where the last equality follows from $\Gamma_{1}=\mathcal{O}(\frac{\kappa_{g}}{\mu}L_{y\theta}^{\Phi})$, $\Gamma_{4}+2\mathbf{C}_{v2}\Gamma_{1}=\mathcal{O}(\kappa_{g}^{3}+L_{y\theta}^{\Phi}\kappa_{g}^{2}/\mu)$, and $\mathbf{L}_{\mathcal{L}_{2}}=\mathcal{O}(\kappa_{g})$.
Next, we focus on the dual gap function in (5.2). Let us define $\bar{K}\triangleq\max\{\frac{2}{1-\rho_{D}}\log(\frac{1}{(1-\beta)\mu}),\frac{2}{1-\beta}\log(\frac{1}{(1-\rho_{D})\mu})\}$. Then, for any $k\geq\bar{K}$, we have that $\max\{\rho_{D}^{k/2}/(1-\beta),\beta^{k/2}/(1-\rho_{D})\}\leq\mu$. Therefore, for any $k\geq\bar{K}$ the RHS in (5.2) can be further bounded as follows
$\mathcal{G}_{\mathcal{Y}}(x_{k},y_{k})$ $\leq\frac{2}{\sigma}\mu^{2}(1-\beta)^{2}\left\|y_{0}-y^{*}_{\mu}(x_{0})\right\|+\frac{2\gamma L_{y\mu}}{\sigma\mu(1-\rho_{D})}D_{\mathcal{X}}+2L_{y\theta}^{\Phi}\beta\mu\left\|\theta_{0}-\theta^{*}(x_{0})\right\|$
$\quad+\frac{2\gamma\beta L_{y\theta}^{\Phi}\mathbf{L}_{\theta}}{(1-\rho_{D})(1-\beta)}D_{\mathcal{X}}+\frac{L_{y\mu}}{\mu\sigma}\gamma D_{\mathcal{X}}+L_{y\theta}^{\Phi}\mu^{2}(1-\rho_{D})^{2}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\frac{\gamma L_{y\theta}^{\Phi}\mathbf{L}_{\theta}}{1-\beta}D_{\mathcal{X}}+\mu D_{\mathcal{Y}}$
$=\mathcal{O}(\frac{\gamma\kappa_{g}}{\mu^{2}}+\frac{\gamma\kappa_{g}^{2}}{\mu}+\mu).$ (71)
where the last equality follows from $1-\rho_{D}=\Omega(\mu)$, $1-\beta=\Omega(1/\kappa_{g})$, and $L_{y\mu}=\mathcal{O}(\kappa_{g})$. Now combining the primal and dual gap functions we conclude that $\mathcal{G}_{\mathcal{Z}}(x_{k^{*}},y_{k^{*}})\leq\mathcal{O}(\frac{\gamma}{\mu^{2}}+\frac{\gamma\kappa_{g}^{2}}{\mu}+\mu+\frac{1}{\sqrt{\gamma\tau K}}+\sqrt{\frac{\kappa_{g}^{3}+\kappa_{g}^{2}/\mu}{\tau K}})$. Selecting $\mu=\mathcal{O}(\kappa_{g}^{3/5}/K^{1/5})$ and $\gamma=\mathcal{O}(\kappa_{g}^{9/5}/K^{3/5})$, one can verify that the primal step-size $\tau$ can be selected as a constant since $\tau\leq\mathcal{O}(\mu^{2}/(\gamma\kappa_{g}^{3}))$ which implies that $\mathcal{G}_{\mathcal{Z}}(x_{k^{*}},y_{k^{*}})\leq\mathcal{O}(\frac{\kappa_{g}^{3/5}}{K^{1/5}}+\frac{\kappa_{g}^{16/5}}{K^{2/5}}+\frac{\kappa_{g}^{3}}{\sqrt{K}})$. Therefore, to achieve $\mathcal{G}_{\mathcal{Z}}(x_{k^{*}},y_{k^{*}})\leq\epsilon$, Algorithm 2 requires at least $K=\mathcal{O}(\frac{\kappa_{g}^{3}}{\epsilon^{5}}+\frac{\kappa_{g}^{8}}{\epsilon^{2.5}}+\frac{\kappa_{g}^{6}}{\epsilon^{2}})$ iterations.
Corollary 5.6.
Suppose the premises of Theorem 5.4 holds and we further assume that $L_{yy}^{\Phi}=0$. Let $\mu=\mathcal{O}(\epsilon)$, $\gamma=\mathcal{O}(\epsilon^{2})$ then for any $K\geq 1$, there exists $t\in\{0,\ldots,K-1\}$ such that $(x_{t},y_{t})\in\mathcal{X}\times\mathcal{Y}$ satisfies $\mathcal{G}_{\mathcal{Z}}(x_{t},y_{t})\leq\epsilon$ within $\mathcal{O}(\frac{1}{\epsilon^{4}}+\frac{\kappa_{g}^{4}}{\epsilon^{2}})$ iterations.
6 Numerical Experiments
In this section, we evaluate the performance of our algorithms in solving the SPB problem. Specifically, we implement our algorithms to address a robust multi-task linear regression example as described in Section 1.3. We utilize both real-world and synthetic datasets, comparing the convergence results of the proposed algorithms across these datasets. For all algorithms, the step-sizes are selected based on the suggestions provided by the theoretical evaluations and are adjusted for enhanced performance across each dataset considered {NLSY79, MTL-dataset, synthetic dataset}. In particular, For i-BRPD:OPF, we let $\gamma=\{0.3,0.7,0.3\}\times\frac{1}{K^{2/3}}$, and $\mu=\{0.1,0.1,0.05\}\times\frac{1}{K^{1/3}}$; for i-BRPD:FP, we let $\gamma=\{1,9,0.05\}\times\frac{1}{K^{1/2}}$, $\mu=\{0.01,0.2,0.01\}\times\frac{1}{K^{1/4}}$, and $\tau=\{0.7,0.7,0.7\}$; in both proposed algorithms, we consider $\alpha=\eta=\frac{2}{\mu_{g}+L_{g}}$, and $\sigma=\frac{2}{2\mu}$; for MORBiT [20], we let $\sigma=\{1,0.01,0.001\}\times\frac{1}{K^{3/5}}$, $\alpha=\{1,1,1\}\times\frac{1}{K^{2/5}}$, $\tau=\{0.1,0.1,0.01\}\times\frac{1}{K^{3/5}}$. In all algorithms, we set the maximum number of iterations $K=10^{4}$.
6.1 Robust Multi-task Linear Regression
In this section, we investigate the performance of our proposed algorithms for solving a robust multi-task learning Problem (1.3). To do this, we consider a robust multi-task linear regression example. In particular, each task $i\in\{1,\ldots,N\}$ consists of a linear regression with training and validation datasets denoted by $\mathcal{D}_{i}^{tr}$ and $\mathcal{D}_{i}^{val}$, respectively. For each task $i$, we consider a linear regression model in the form of $b_{i}=A_{i}(\lambda_{i}y_{i}+(1-\lambda_{i})x)+\epsilon_{i}$, which is a linear combination of the common coefficients $(x)$ and task-specific coefficients $(y_{i})$ using a given dataset $\mathcal{D}_{i}=(A_{i},b_{i})$, where $A_{i}\in\mathbb{R}^{n_{i}\times d}$ and $b_{i}\in\mathbb{R}^{n_{i}}$. We define the loss function as $\ell_{i}(y_{i};\mathcal{D}_{i})=\frac{1}{2n_{i}}\left\|A_{i}y_{i}-b_{i}\right\|^{2}$ and assume the feature space is shared among all tasks. This problem can be formulated as follows:
$\min_{\|x\|_{1}\leq Q,\lambda\in[0,1]^{N}}\max_{i=1}^{N}~{}\ell_{i}(y_{i}^{*}(x,\lambda_{i});\mathcal{D}_{i}^{val})\equiv\min_{\|x\|_{1}\leq Q,\lambda\in[0,1]^{N}}\max_{\eta\in\Delta_{N}}\sum_{i=1}^{N}\eta_{i}\ell_{i}(y_{i}^{*}(x,\lambda_{i});\mathcal{D}_{i}^{val})$
[…]
where we encounter a strongly convex lower-level (LL) problem, while the upper-level (UL) problem is a nonconvex-concave saddle point (SP) problem similar to our main Problem (1).
To showcase the performance of the proposed algorithms in solving Problem (6.1) across various datasets, we examine three regression-based datasets. Firstly, we employ the 1979 National Longitudinal Survey of Youth (NLSY79) dataset [^1] NLSY79 dataset overview ($n=6213,d=21$). Next, we utilize the MTL-dataset ($n=5642,d=14$), which comprises a mixture of datasets including bodyfat, housing, mg, mpg, and space-ga from the LIBSVM library [7]. Additionally, we include a synthetic dataset ($n=5000,d=100$), generated randomly from a standard Normal distribution. For these datasets, we conduct experiments with $N=5$ tasks. We distribute the data points evenly among the considered tasks. Then, we allocate 75% of the data for each task as the training dataset and reserve the remaining 25% as the validation dataset. To evaluate the performance of our proposed algorithms (i-BRPD:OPF and i-BRPD:FP), we compared them with the Multi-Objective Robust Bilevel Two-timescale optimization (MORBiT) algorithm introduced by [20] using three distinct datasets. As depicted in Figure 1, our proposed algorithms exhibit a faster convergence rate compared to the MORBiT algorithm, aligning with our theoretical findings. Specifically, the i-BRPD:OPF algorithm, which is a one-sided projection-free method, demonstrates the fastest convergence behavior among all algorithms. This can be attributed to the utilization of the Linear Minimization Oracle (LMO) for updating the primal variable of the objective function, circumventing the need for complex projections at each iteration as required by projection-based methods. Moreover, our fully projected gradient-based approach, i-BRPD:FP, surpasses the projection-based MORBiT algorithm. This superiority stems from our proposed algorithm’s utilization of an inexact approach for estimating the Hessian inverse matrix at each iteration, as discussed in Section 3. In contrast, the MORBiT method approximates the Hessian inverse using the Neumann series, as discussed in [15], which is less accurate and computationally expensive compared to our approach.
(a) NLSY97 dataset
(b) MTL-dataset
(c) Synthetic dataset
Figure 1: Comparing the performance of our proposed algorithms i-BRPD:OPF (blue) and i-BRPD:FP (red) with MORBiT (green) in Robust Multi-task Linear Regression problem
7 Conclusions
In this paper, we introduced a class of constrained saddle point problems with a bilevel structure (SPB), where the upper-level objective function is nonconvex-concave and smooth over a compact and convex constraint set, subject to a strongly convex lower-level objective function. By leveraging regularization and nested approximation techniques, we initially devised a bilevel primal-dual one-sided projection-free algorithm (i-BRPD:OPF), requiring $\mathcal{O}(\epsilon^{-4})$ iterations to attain an $\epsilon$-stationary point. Subsequently, we developed a bilevel primal-dual fully projected algorithm (i-BRPD:FP), capable of achieving an $\epsilon$-stationary solution within $\mathcal{O}(\epsilon^{-5})$ iterations. Our numerical experiments revealed the superior performance of our proposed algorithms compared to existing methods, which predominantly rely on projection-based approaches to address nonconvex strongly-concave upper-level objective functions or upper-level objective functions that are linear in the dual variable.
Acknowledgements
This work is supported in part by the Arizona Technology and Research Initiative Fund (TRIF) for Innovative Technologies for the Fourth Industrial Revolution.
Appendix
In this section, we introduce a collection of technical lemmas necessary to support the results presented in the paper.
8 Required Lemmas
In this section, we provide detailed explanations and proofs for the lemmas supporting the main results of the paper.
Lemma 8.1.
Let $\{A_{k}\}_{k\geq 0}$ be a sequence of nonnegative real numbers, $\rho\in(0,1)$, and $\beta\in(0,1)$. For any $K\geq 1$,
  • a)
    $\sum_{k=0}^{K-1}\sum_{i=0}^{k}\beta^{k-i}A_{i}\leq\frac{1}{1-\beta}\sum_{k=0}^{K-1}A_{k}$.
  • b)
    $\sum_{k=0}^{K-1}A_{k}\sum_{i=0}^{k-1}\beta^{k-i}A_{i}\leq\frac{\beta}{1-\beta}\sum_{k=0}^{K-1}A_{k}^{2}$.
  • c)
    $\sum_{k=0}^{K-1}A_{k}\sum_{i=0}^{k}\rho^{k-i}\sum_{j=0}^{i-1}\beta^{i-j}A_{j}\leq\frac{\beta}{(1-\beta)(1-\rho)}\sum_{k=0}^{K-1}A_{k}^{2}$.
Proof.
Part (a):
$\sum_{k=0}^{K-1}\sum_{i=0}^{k}\beta^{k-i}A_{i}$ $=A_{0}\sum_{k=0}^{K-1}\beta^{i}+\sum_{k=0}^{K-2}\beta^{i}+\sum_{k=0}^{K-3}\beta^{i}+...+\sum_{k=0}^{K-1-(K-1)}\beta^{i}$
$\leq\frac{1}{1-\beta}(A_{0}+A_{1}+...+A_{K-1})$
$\leq\frac{1}{1-\beta}\sum_{k=0}^{K-1}A_{k}.$ (73)
Part (b): Using Young’s inequality and part (a) we can show
$\sum_{k=0}^{K-1}A_{k}\sum_{i=0}^{k-1}\beta^{k-i}A_{i}$ $\leq\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\beta^{k-i}(A_{k}^{2}+A_{i}^{2})/2$
$\leq\sum_{k=0}^{K-1}\Big{(}\frac{\beta}{1-\beta}\frac{A_{k}^{2}}{2}+\frac{\beta}{1-\beta}\frac{A_{k}^{2}}{2}\Big{)}$
$\leq\frac{\beta}{1-\beta}\sum_{k=0}^{K-1}A_{k}^{2}.$ (74)
To prove part (c), first using Young’s inequality we have that
$\sum_{k=0}^{K-1}A_{k}\sum_{i=0}^{k}\rho^{k-i}\sum_{j=0}^{i-1}\beta^{i-j}A_{j}$ $\leq\sum_{k=0}^{K-1}\sum_{i=0}^{k}\sum_{j=0}^{i-1}\rho^{k-i}\beta^{i-j}(A_{k}^{2}+A_{j}^{2})/2$
$\leq\sum_{k=0}^{K-1}\left(\frac{\beta}{(1-\beta)(1-\rho)}\frac{A_{k}^{2}}{2}+\sum_{i=0}^{k}\sum_{j=0}^{i-1}\rho^{k-i}\beta^{i-j}\frac{A_{j}^{2}}{2}\right).$ (75)
Now, let us define $A^{\prime}_{i}\triangleq\sum_{j=0}^{i-1}\beta^{i-j}A_{j}^{2}$, then applying part (a) twice we obtain $\sum_{k=0}^{K-1}\sum_{i=0}^{k}\rho^{k-i}A^{\prime}_{i}\leq\frac{1}{(1-\rho)}\sum_{k=0}^{K-1}A^{\prime}_{k}=\frac{1}{(1-\rho)}\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\beta^{k-i}A_{i}^{2}\leq\frac{\beta}{(1-\rho)(1-\beta)}\sum_{k=0}^{K-1}A_{k}^{2}$. Therefore, using this inequality within (8) the desired result can be concluded.
8.1 Proof of Lemma 2.1
Before presenting the proof of Lemma 2.2, we show a preliminary result essential in establishing the result of Lemma 2.2. In particular, we show that the map $v(x,y)$ defined in (8b) is Lipschitz continuous.
Proof.
We start the proof by recalling that $v(x,y)=[\nabla_{\theta\theta}^{2}g(x,\theta^{*}(x))]^{-1}\nabla_{\theta}\Phi(x,\theta^{*}(x),y)$. Next, adding and subtracting $[\nabla_{\theta\theta}^{2}g(x,\theta^{*}(x))]^{-1}\nabla_{\theta}\Phi(\overline{x},\theta^{*}(\overline{x}),\overline{y})$ followed by a triangle inequality leads to,
$\left\|v(x,y)-v(\overline{x},\overline{y})\right\|$ $=\left\|[\nabla_{\theta\theta}^{2}g(x,\theta^{*}(x))]^{-1}\nabla_{\theta}\Phi(x,\theta^{*}(x),y)\right.$
$\quad-\left.[\nabla_{\theta\theta}^{2}g(\overline{x},\theta^{*}(\overline{x}))]^{-1}\nabla_{\theta}\Phi(\overline{x},\theta^{*}(\overline{x}),\overline{y})\right\|$
$\leq\left\|[\nabla_{\theta\theta}^{2}g(x,\theta^{*}(x))]^{-1}(\nabla_{\theta}\Phi(x,\theta^{*}(x),y)-\nabla_{\theta}\Phi(\overline{x},\theta^{*}(\overline{x}),\overline{y}))\right\|$
$\quad+\left\|([\nabla_{\theta\theta}^{2}g(x,\theta^{*}(x))]^{-1}-[\nabla_{\theta\theta}^{2}g(\overline{x},\theta^{*}(\overline{x}))]^{-1})\nabla_{\theta}\Phi(\overline{x},\theta^{*}(\overline{x}),\overline{y})\right\|$
$\leq\frac{1}{\mu_{g}}(L_{\theta x}^{\Phi}\left\|x-\overline{x}\right\|+L_{\theta\theta}^{\Phi}\left\|\theta^{*}(x)-\theta^{*}(\overline{x})\right\|+L_{\theta y}^{\Phi}\left\|\overline{y}-y\right\|)$
$\quad+C_{\theta}^{\Phi}\left\|[\nabla_{\theta\theta}^{2}g(x,\theta^{*}(x))]^{-1}-[\nabla_{\theta\theta}^{2}g(\overline{x},\theta^{*}(\overline{x}))]^{-1}\right\|.$ (76)
where in the last inequality we used Assumptions 2.1 and 2.2-(iii) along with the premises of Assumption 2.3. Moreover, for any invertible matrices $H_{1}$ and $H_{2}$, we have that
$\left\|H_{2}^{-1}-H_{1}^{-1}\right\|=\left\|H_{1}^{-1}(H_{1}-H_{2})H_{2}^{-1}\right\|\leq\left\|H_{1}^{-1}\right\|\left\|H_{2}^{-1}\right\|\left\|H_{1}-H_{2}\right\|.$ (77)
Therefore, using the result of Lemma 2.2-(I) and (77) we can further bound inequality (8.1) as follows,
$\left\|v(x,y)-v(\overline{x},\overline{y})\right\|$ $\leq\frac{1}{\mu_{g}}(L_{\theta x}^{\Phi}\left\|x-\overline{x}\right\|+L_{\theta\theta}^{\Phi}\mathbf{L}_{\theta}\left\|x-\overline{x}\right\|+L_{\theta y}^{\Phi}\left\|\overline{y}-y\right\|)$
$\quad+C_{\theta}^{\Phi}\left\|[\nabla_{\theta\theta}^{2}g(x,\theta^{*}(x))]^{-1}-[\nabla_{\theta\theta}^{2}g(\overline{x},\theta^{*}(\overline{x}))]^{-1}\right\|$
$\leq\frac{1}{\mu_{g}}(L_{\theta x}^{\Phi}+L_{\theta\theta}^{\Phi}\mathbf{L}_{\theta})\left\|x-\overline{x}\right\|+\frac{1}{\mu_{g}}L_{\theta y}^{\Phi}\left\|\overline{y}-y\right\|$
$\quad+\frac{C_{\theta}^{\Phi}}{\mu^{2}_{g}}L_{\theta\theta}^{g}(\left\|x-\overline{x}\right\|+\left\|\theta^{*}(x)-\theta^{*}(\overline{x})\right\|)$
$\leq(\frac{L_{\theta x}^{\Phi}+L_{\theta\theta}^{\Phi}\mathbf{L}_{\theta}}{\mu_{g}}+\frac{C_{\theta}^{\Phi}L_{\theta\theta}^{g}}{\mu^{2}_{g}}(1+\mathbf{L}_{\theta}))\left\|x-\overline{x}\right\|+\frac{1}{\mu_{g}}L_{\theta y}^{\Phi}\left\|\overline{y}-y\right\|.$ (78)
The result follows from the definitions of $\mathbf{C}_{v1}$ and $\mathbf{C}_{v2}$.
8.2 Proof of Lemma 2.2
Part (I): Recall that $\theta^{*}(x)$ is the minimizer of the lower-level problem whose objective function is strongly convex, therefore
$\mu_{g}\left\|\theta^{*}(x)-\theta^{*}(\overline{x})\right\|^{2}$ $\leq\langle\nabla_{\theta}g(x,\theta^{*}(x))-\nabla_{\theta}g(x,\theta^{*}(\overline{x})),\theta^{*}(x)-\theta^{*}(\overline{x})\rangle$
$=\langle\nabla_{\theta}g(\overline{x},\theta^{*}(\overline{x}))-\nabla_{\theta}g(x,\theta^{*}(\overline{x})),\theta^{*}(x)-\theta^{*}(\overline{x})\rangle.$
Note that $\nabla_{\theta}g(x,\theta^{*}(x))=\nabla_{\theta}g(\overline{x},\theta^{*}(\overline{x}))=0$. Using the Cauchy-Schwartz inequality we have
$\mu_{g}\left\|\theta^{*}(x)-\theta^{*}(\overline{x})\right\|^{2}$ $\leq\left\|\nabla_{\theta}g(\overline{x},\theta^{*}(\overline{x}))-\nabla_{\theta}g(x,\theta^{*}(\overline{x}))\right\|\left\|\theta^{*}(x)-\theta^{*}(\overline{x})\right\|$
$\leq C_{\theta x}^{g}\left\|x-\overline{x}\right\|\left\|\theta^{*}(x)-\theta^{*}(\overline{x})\right\|.$
where the last inequality is obtained by using Assumption 2.2. Therefore, we conclude that $\mu_{g}\left\|\theta^{*}(x)-\theta^{*}(\overline{x})\right\|\leq C_{\theta x}^{g}\left\|x-\overline{x}\right\|$ which leads to the desired result in part (I).
Part (II): We start proving this part using the definition of $\nabla_{x}\mathcal{L}(x,y)$ stated in (8a). Using the triangle inequality we obtain,
$\left\|\nabla_{x}\mathcal{L}(x,y)-\nabla_{x}\mathcal{L}(\overline{x},\overline{y})\right\|$ $=\left\|\nabla_{x}\Phi(x,\theta^{*}(x),y)-\nabla_{\theta x}^{2}g(x,\theta^{*}(x))v(x,y)\right.$
$\quad-\left.(\nabla_{x}\Phi(\overline{x},\theta^{*}(\overline{x}),\overline{y})-\nabla_{\theta x}^{2}g(\overline{x},\theta^{*}(\overline{x}))v(\overline{x},\overline{y})\right\|$
$\leq\left\|\nabla_{x}\Phi(x,\theta^{*}(x),y)-\nabla_{x}\Phi(\overline{x},\theta^{*}(\overline{x}),\overline{y})\right\|$
$\quad+\left\|[\nabla_{\theta x}^{2}g(\overline{x},\theta^{*}(\overline{x}))v(\overline{x},\overline{y})-\nabla_{\theta x}^{2}g(\overline{x},\theta^{*}(\overline{x}))v(x,y)]\right.$
$\quad+\left.[\nabla_{\theta x}^{2}g(\overline{x},\theta^{*}(\overline{x}))v(x,y)-\nabla_{\theta x}^{2}g(x,\theta^{*}(x))v(x,y)]\right\|,$ (79)
where the second term of the RHS follows from adding and subtracting the term $\nabla_{\theta x}^{2}g(\overline{x},\theta^{*}(\overline{x}))v(x,y)$.
Next, from Assumptions 2.1-(1) and 2.2-(ii) together with the triangle inequality we conclude that
$\left\|\nabla_{x}\mathcal{L}(x,y)-\nabla_{x}\mathcal{L}(\overline{x},\overline{y})\right\|$ $\leq L_{xx}^{\Phi}\left\|x-\overline{x}\right\|+L_{x\theta}^{\Phi}\left\|\theta^{*}(x)-\theta^{*}(\overline{x})\right\|+L_{xy}^{\Phi}\left\|\overline{y}-y\right\|$
$\quad+C_{\theta x}^{g}\left\|v(\overline{x},\overline{y})-v(x,y)\right\|+\frac{C_{\theta}^{\Phi}}{\mu_{g}}\left\|\nabla_{\theta x}^{2}g(\overline{x},\theta^{*}(\overline{x}))-\nabla_{\theta x}^{2}g(x,\theta^{*}(x))\right\|$ (80)
Note that in the last inequality, we use the fact that $\left\|v(x,y)\right\|=\left\|[\nabla_{\theta\theta}^{2}g(x,\theta^{*}(x))]^{-1}\nabla_{\theta}\Phi(x,\theta^{*}(x),y)\right\|\leq\frac{C_{\theta}^{\Phi}}{\mu_{g}}$.
Combining the result of Lemma 2.2-(I) and Assumption 2.3 with the Assumption 2.2-(iv) leads to
$\left\|\nabla_{x}\mathcal{L}(x,y)-\nabla_{x}\mathcal{L}(\overline{x},\overline{y})\right\|$ $\leq L_{xx}^{\Phi}\left\|x-\overline{x}\right\|+L_{x\theta}^{\Phi}\mathbf{L}_{\theta}\left\|x-\overline{x}\right\|+L_{xy}^{\Phi}\left\|\overline{y}-y\right\|$
$\quad+C_{\theta x}^{g}(\mathbf{C}_{v1}\left\|x-\overline{x}\right\|+\mathbf{C}_{v2}\left\|y-\overline{y}\right\|)$
$\quad+\frac{C_{\theta}^{\Phi}}{\mu_{g}}L_{\theta x}^{g}(\left\|\ x-\overline{x}\right\|+\left\|\ \theta^{*}(x)-\theta^{*}(\overline{x})\right\|)$
$\leq L_{xx}^{\Phi}\left\|x-\overline{x}\right\|+L_{x\theta}^{\Phi}\mathbf{L}_{\theta}\left\|x-\overline{x}\right\|+L_{xy}^{\Phi}\left\|\overline{y}-y\right\|$
$\quad+C_{\theta x}^{g}(\mathbf{C}_{v1}\left\|x-\overline{x}\right\|+\mathbf{C}_{v2}\left\|y-\overline{y}\right\|)$
$\quad+\frac{C_{\theta}^{\Phi}}{\mu_{g}}L_{\theta x}^{g}(\left\|\ x-\overline{x}\right\|+\mathbf{L}_{\theta}\left\|x-\overline{x}\right\|)$
$\leq\Big{(}L_{xx}^{\Phi}+L_{x\theta}^{\Phi}\mathbf{L}_{\theta}+C_{\theta x}^{g}\mathbf{C}_{v1}+\frac{C_{\theta}^{\Phi}}{\mu_{g}}L_{\theta x}^{g}(1+\mathbf{L}_{\theta})\Big{)}\left\|x-\overline{x}\right\|$
$\quad+(L_{xy}^{\Phi}+C_{\theta x}^{g}\mathbf{C}_{v2})\left\|y-\overline{y}\right\|,$ (81)
where $\mathbf{L}_{\mathcal{L}_{1}}=(L_{xx}^{\Phi}+L_{x\theta}^{\Phi}\mathbf{L}_{\theta}+C_{\theta x}^{g}\mathbf{C}_{v1}+\frac{C_{\theta}^{\Phi}}{\mu_{g}}L_{\theta x}^{g}(1+\mathbf{L}_{\theta}))$ and $\mathbf{L}_{\mathcal{L}_{2}}=(L_{xy}^{\Phi}+C_{\theta x}^{g}\mathbf{C}_{v2})$.
Part (III): Using the definition of $\nabla_{y}\mathcal{L}(x,y)$ in (6b) and Assumption 2.1-(3), we can show
$\|\nabla_{y}\mathcal{L}(x,y)-\nabla_{y}\mathcal{L}(\overline{x},\overline{y})\|$ $=\|\nabla_{y}\Phi(x,\theta^{*}(x),y)-\nabla_{y}\Phi(\overline{x},\theta^{*}(x),\overline{y})\|$
$\leq L_{yx}^{\Phi}\left\|x-\overline{x}\right\|+L_{y\theta}^{\Phi}\|\theta^{*}(x)-\theta^{*}(\overline{x})\|+L_{yy}^{\Phi}\left\|y-\overline{y}\right\|$
$\leq L_{yx}^{\Phi}\left\|x-\overline{x}\right\|+L_{y\theta}^{\Phi}\mathbf{L}_{\theta}\left\|x-\overline{x}\right\|+L_{yy}^{\Phi}\left\|y-\overline{y}\right\|$
$\leq(L_{yx}^{\Phi}+L_{y\theta}^{\Phi}\mathbf{L}_{\theta})\left\|x-\overline{x}\right\|+L_{yy}^{\Phi}\left\|y-\overline{y}\right\|.$ (82)
where the last inequality is obtained by using Lemma 2.2-(I).